### v1

Abstract
We introduce OneCAT, a unified multimodal model that seamlessly integrates understanding, generation, and editing within a novel, pure decoder-only transformer architecture. Our framework uniquely eliminates the need for external components such as Vision Transformers (ViT) or vision tokenizer during inference, leading to significant efficiency gains, especially for high-resolution inputs. This is achieved through a modality-specific Mixture-of-Experts (MoE) structure trained with a single autoregressive (AR) objective, which also natively supports dynamic resolutions. Furthermore, we pioneer a multi-scale visual autoregressive mechanism within the Large Language Model (LLM) that drastically reduces decoding steps compared to diffusion-based methods while maintaining state-of-the-art performance. Our findings demonstrate the powerful potential of pure autoregressive modeling as a sufficient and elegant foundation for unified multimodal intelligence. As a result, OneCAT sets a new performance standard, outperforming existing open-source unified multimodal models across benchmarks for multimodal generation, editing, and understanding.
1 Introduction
In the past few years, modular approaches, utilizing separate architectures for understanding [2, 65, 3, 13, 14, 12], generation [40, 36, 59, 24, 34, 20], and editing tasks [85, 5, 32, 87, 46, 86], dominated multi-modal frameworks. While producing capable systems, it inherently creates complex, multi-stage pipelines. Such designs often suffer from architectural bottlenecks that limit deep, early-stage fusion of cross-modal information and introduce significant inference latency, presenting a major barrier to both efficiency and performance. In response to these limitations, the field has seen a rapid convergence towards unified multimodal modeling, aiming to integrate these disparate capabilities within a single, end-to-end architecture [67, 62, 75, 76, 11, 16]. Despite the trend towards unification, many current models remain tethered to the modular paradigm. We contend that a true paradigm shift—necessary to unlock the full potential of unified systems—demands a move towards a more fundamental, first-principles architecture that eschews external components. We therefore propose that a pure decoder-only transformer, trained under a unified objective, provides not just a sufficient, but a more elegant and potent foundation for the next generation of general-purpose multimodal intelligence.
Figure 1: Showcase of the text-to-image generation abilities of the OneCAT model.
Figure 2: Showcase of the image editing abilities of the OneCAT model, including general image editing tasks such as object removal, background adjustment, color adjustment, subject replacement, and style transfer; as well as perceptual tasks including depth estimation, pose estimation, object segmentation, and Canny edge detection.
To realize this vision, we re-evaluate the core architectural tenets of multimodal systems. We introduce an encoder-free framework where raw visual inputs are directly tokenized into patch embeddings and processed alongside text tokens within a single decoder stack. The critical innovation is a modality-specific Mixture-of-Experts (MoE) layer, which dynamically routes vision and text tokens to specialized experts, enabling deep, early-stage feature fusion in an efficient inference manner without the need for exquisite encoders. For generative tasks, we pioneer a multi-scale autoregressive mechanism [24, 63] inside the LLM, augmented with the proposed scale-aware adapter modules, to predict image tokens from low to high resolution. This design not only circumvents the high latency of iterative diffusion models but also allows the model to learn a coarse-to-fine generative process, significantly enhancing both speed and output quality.
Regarding data protocol, we leverage mixed training strategies that combine large-scale, web-scraped image-text pairs with curated, instruction-following datasets. Our pre-training corpus is constructed to expose the model to a diverse range of data formats, including image-text documents, visual question-answering pairs, and generative prompts for both text-to-image synthesis and image editing. This heterogeneous data mixture is crucial for training a truly unified model, as it forces the shared decoder to develop a generalized representation that can seamlessly switch between comprehension, generation, and editing tasks. We argue that this unified data diet, coupled with our bottleneck-free architecture, is key to fostering the synergistic emergence of complex multimodal capabilities.
Building upon these principles, we present Only DeCoder Auto-regressive Transformer (OneCAT), an open-source unified multimodal model. Comprehensive evaluations demonstrate that OneCAT sets a new state-of-the-art for pure decoder-only unified models. More importantly, the encoder-free design provides a significant inference speedup, particularly for high-resolution inputs. By demonstrating the viability and superiority of a pure decoder architecture, OneCAT offers a more first-principles-aligned paradigm for multimodal modeling. It facilitates earlier cross-modal fusion through its unified MoE structure and enhances semantic consistency, providing valuable insights and a powerful new baseline for the development of next-generation unified multimodal systems. For examples of our model’s impressive image generation capabilities, please refer to Fig. 1. Its advanced image editing functionalities are further showcased in Fig. 2.
2 Related Work
2.1 Compositional MLLMs
The field of Multimodal Large Language Models (MLLMs) has rapidly evolved, converging on a dominant compositional architecture. This paradigm connects a pre-trained vision encoder, such as CLIP [56], SigLIP [84], or the more recent InternViT [14], to a powerful Large Language Model (LLM) through a trainable connector. Pioneering works [1, 37] propose sophisticated connector designs. For example, Flamingo [1] introduces gated cross-attention layers to inject visual information into an LLM, while BLIP-2 [37] develops the Q-Former to bridge the modality gap between an image encoder and an LLM. A significant shift occurs with LLaVA [45], which simplifies the connector to a lightweight Multi-Layer Perceptron (MLP) projection layer. This effective design become a foundational blueprint for numerous subsequent MLLMs. For example, recent state-of-the-art models like the InternVL series [14, 12, 13, 91, 66] and the Qwen-VL series [65, 3, 2] adopt this same core principle and achieve superior performance by leveraging larger-scale training data and more powerful vision and language foundation models. However, this successful compositional design has inherent drawbacks. The separate nature of the vision and language components complicates the end-to-end optimization process and introduces two critical bottlenecks. First, the sequential nature of the architecture, where the vision encoder must fully process an image before the LLM can begin its generation, leads to high inference latency. This initial step is often referred to as the “prefilling” stage. Second, the connector itself acts as an information bottleneck. In this so-called late fusion pipeline, complex visual information is compressed into a compact representation for the LLM, inevitably causing a loss of fine-grained visual detail. These fundamental limitations are now motivating a shift in the field towards more deeply integrated, such as decoder-only models, that aim to overcome these challenges.
2.2 Decoder-only MLLMs
Decoder-only MLLMs, also known as monolithic MLLMs, have recently emerged as a minimalist yet powerful alternative to the conventional compositional architecture. This paradigm aims to achieve greater efficiency and a more direct early fusion of modalities by eliminating the separate vision encoder. For example, Fuyu-8B [4] processes vision patches by feeding them through a simple linear patch embedding layer directly into the LLM backbone, which markedly reduce inference latency. Inspired by this success, subsequent works [18, 49, 19] further advance decode-only MLLMs by targeting their training processes and architectures. EvE [18] introduces a novel training objective that guides the model’s visual learning by aligning the last layer of LLM’s hidden states of image patch tokens with semantic features from a pre-trained vision encoder. Differently, Mono-InternVL [49] and EvEv2.0 [19] adopt a Mixture-of-Experts (MoE) framework, introducing a dedicated visual expert to handle visual-specific features more effectively. HoLVE [61] prepends a causal transformer to the LLM to explicitly convert both visual and textual inputs into a shared space. Despite these promising advancements, the overall training efficiency of these models remains a significant challenge. More importantly, the potential for the decoder-only architecture to create unified models that can seamlessly integrate multimodal understanding, generation, and even image editing capabilities remains a largely unexplored research avenue.
2.3 Unified Visual Understanding and Generation
Building on the success of MLLMs, the convergence of visual understanding and generation into a unified framework now represents a key research frontier. Pioneering unified MLLMs such as Chameleon [62],Transfusion [90], emu3 [67] and show-o [75] utilize visual tokenizers (e.g., VQ-VAE) to convert images into a sequence of discrete tokens, enabling seamless multimodal understanding and generation within a single model. However, the discretization inevitably results in lossy visual information and weakens in extracting semantic contents. Janus series [69, 11] decouples visual encoding for understanding and generation using two separate encoders, but may compromise performance due to task conflicts in shared LLM parameter space. Metaqueries [53], BLIP3-O [8], Uniworld-V [43] assembles off-the-shelf specialized MLMMs and diffusion models by tuning adapters and learnable query tokens, which sacrifices true architectural unification for modularity. BAGEL [16] and Mogao [41] employ a Mixture-of-Transformers (MoT) architecture, dedicating different components to autoregressive text generation and diffusion-based visual generation. However, while powerful, this hybrid approach inherits the significant inference latency of diffusion models and still requires separate encoders and tokenizers during the inference.
In contrast to these approaches, our OneCAT introduces a pure decoder-only architecture. By integrating modality-specific experts directly within the decoder, OneCAT achieves versatile multimodal capabilities without the need for external vision encoders or tokenizers at inference time, thus resolving the trade-off between architectural purity and inference efficiency.
2.4 Next Scale Prediction for Visual Generation
Autoregressive models based on next-token prediction(NTP) have long faced efficiency challenges in high-resolution image generation due to the quadratic growth of sequence length with image size. Similarly, diffusion models—though widely successful—often suffer from slow iterative sampling. To address these limitations, VAR [63] introduced the next-scale prediction(NSP) paradigm, which encodes images into hierarchical discrete tokens via a multi-scale VAE and generates them autoregressively from low to high resolution, significantly reducing the number of decoding steps. Building upon this, Infinity [24] further enhanced this approach with bit-level prediction and extended tokenizer vocabulary, achieving superior generation quality while maintaining efficient inference. To enable unified understanding and generation, VARGPT [92] stack the transformer from pretrained VAR [63] as a visual decoder atop a LLM. However, since the visual tokens (i.e., the input of the visual decoder) must be decoded token-by-token through the LLM before subsequent next-scale prediction, this approach compromises the inference efficiency that is the key advantage of the NSP.
In contrast, our proposed OneCAT seamlessly unifies next-token prediction for text generation and next-scale prediction for visual generation within a single LLM.
3 OneCAT
Figure 3: Overview inference pipeline of OneCAT. OneCAT is a decoder-only autoregressive unified model that seamlessly supports multimodal understanding, text-to-image generation and image editing. OneCAT implements both next-token prediction for text generation and next-scale prediction for visual generation within a unified LLM backbone. In the left part, the dash squares denote the generated text and visual token during inference. In the right part, PE and $T$ denote patch embedding and text tokenizer, respectively. TD and ID denote text detokenizer and image detokenizer, respectively.
As illustrated in Fig. 3, OneCAT employs a pure decoder-only architecture, eliminating the need for any additional vision encoder or tokenizer during inference. This streamlined design significantly simplifies the model structure and reduces computational overhead. Unlike traditional multimodal models that rely on semantic encoders like ViTs for understanding [3, 16, 41, 11], OneCAT utilizes a lightweight Patch Embedding layer. This layer losslessly converts raw images into continuous visual tokens, enabling efficient multimodal understanding. Crucially, the same Patch Embedding layer also encodes reference images for editing tasks, thereby superseding separate VAE-based tokenizers traditionally used [16, 41] and further enhancing inference efficiency.
At its core, OneCAT integrates a Mixture-of-Experts (MoE) architecture. This MoE comprises three specialized feed-forward network (FFN) experts: one dedicated to processing text tokens for language comprehension, another designed for continuous visual tokens to facilitate multimodal understanding, and a third optimized for generating discrete visual tokens during image synthesis. All QKV and attention layers are shared across modalities and tasks, which promotes significant parameter efficiency and robust cross-modal alignment for instruction-following.
For text generation, OneCAT adheres to the conventional Next-Token Prediction paradigm, leveraging the well-established capabilities of autoregressive language modeling. In parallel, for visual generation, it innovatively employs the Next-Scale Prediction paradigm [63, 24]. This mechanism generates images in a coarse-to-fine, hierarchical manner, progressively predicting visual tokens from the lowest to the highest resolution scale, thereby achieving high-quality visual outputs.
3.1 Architecture
Our OneCAT model is initialized from the pre-trained Qwen2.5 LLM [77], leveraging its strong language modeling capabilities as a foundation. To construct our MoE architecture, we replicate the original FFN layer from each Qwen2.5 transformer block to form three distinct experts: a Text FFN (i.e., Text. FFN), a Visual Understanding FFN (i.e., Und. FFN), and a Visual Generation FFN (i.e., Gen. FFN). OneCAT employs a deterministic hard routing mechanism, where tokens are assigned to a specific expert based on their modality and the task at hand. The core functionality for each task is handled as follows:
Multimodal Understanding.
To process images for understanding tasks, we employ a simple yet effective patch embedding layer that converts raw images into a sequence of continuous visual tokens. This layer consists of a $\times14$ convolution for image patchifying, a $\times2$ pixel unshuffle operation for visual token compression, and a two-layer Multilayer Perceptron (MLP) for projecting the visual features to match the LLM’s hidden state dimension. These continuous visual tokens are exclusively routed to the Und. FFN, while the text tokens for instruction are routed to the Text FFN.
Text-to-Image Generation.
For image generation, we leverage a pre-trained multi-scale VAE model from Infinity [24] to map images between pixel space and latent space. This VAE operates with a downsampling ratio of 16 and a latent channel size of 32, and incorporates a bitwise quantizer [89] to enlarge the vocabulary. During training, the image tokenizer transforms the target images into a sequence of multi-scale discrete visual tokens to serve as ground-truth and input of LLM for teacher-forcing training, which are processed by the Gen. FFN. Critically, during inference, the tokenizer is not required; only the detokenizer is needed to reconstruct the final image from the generated multi-scale visual tokens. The conditional text tokens are also routed to the Text FFN.
Image Editing.
OneCAT seamlessly supports image editing task by conditioning the visual generation process on a reference image and edit-instruction. The reference image is processed through the patch embedding layer, and the resulting continuous visual tokens are routed to the Und. FFN to serve as the visual condition, while the text tokens for instruction are routed to the Text FFN. The patch embedding layer provides a near-lossless representation of the reference image, which allows the LLM’s shallower layers to access low-level features for pixel-level consistency, while the deeper layers extract high-level features for semantical comprehension. Guided by this rich, hierarchical visual context, the model then autoregressively predicts new discrete visual tokens, which are handled by the Gen. FFN. This design enables powerful conditional generation without requiring any architectural modifications, showcasing the versatility of our unified decoder-only design.
3.2 Scale-Aware Adapter for Hierarchical Generation
While the Gen. FFN is capable of processing discrete visual tokens, the tokens produced by the multi-scale VAE are inherently hierarchical. A standard FFN would treat tokens for different scales equally, ignoring this crucial structural information. To address this and enable more granular control over the visual generation process, we introduce the Scale-Aware Adapter (SAA), a novel architectural component integrated with the Gen. FFN.
Our design is motivated by the principle that different scales in the multi-scale VAE govern distinct aspects for image generation. Specifically, tokens from lower scales predominantly encode low-frequency global information, such as color, illumination, and coarse structure. Those from higher scales, conversely, capture high-frequency details including fine textures and intricate patterns. Processing these functionally divergent tokens with a shared Gen. FFN layer limits representational capacity and is thus suboptimal.
The Scale-Aware Adapter (SAA) comprises a set of parallel modules that serve as skip connections over the Gen. FFN. Each module is dedicated to processing tokens from a specific scale of the multi-scale VAE, with the total number of adapters matching the number of VAE scales. During inference, discrete visual tokens are routed to their corresponding scale-specific adapter based on the scale index. To ensure parameter efficiency, each adapter is constructed using a low-rank decomposition (rank rr= 64), inspired by the LoRA [26]. However, unlike LoRA which is typically used for fine-tuning, the SAA modules are trained jointly end-to-end as permanent components of the LLM.
3.3 Multimodal versatile attention mechanism
We leverage a multimodal versatile attention mechanism based on PyTorch FlexAttention [55] to empower OneCAT with the ability to process diverse modalities and tasks in a flexible and adaptive manner. As illustrated in Fig. 4, text tokens $T$ are processed using causal attention, ensuring autoregressive generation; Continuous visual tokens $U$ are processed via full attention, allowing each token to attend to all others in the sequence. Multiscale discrete visual tokens $G_{i}$ (where $i$ denotes the scale index) are processed via block causal attention, tokens within the same scale can attend to each other freely, while attention across scales follows a causal attention.
Figure 4: Multimodal versatile attention mechanism. $T$ denotes the text tokens. $U$ denotes the continuous visual tokens for multimodal understanding or reference image tokens for image editing. $G_{i}$ denotes the $i$-th scale discrete visual tokens for visual generation.
4 Model Training Pipeline
The training pipeline is divided into three stages and the instruction of the training recipe is shown in Tab. 1.
Figure 5: Overview of the training pipeline. In Stage 1, we first prepare a teacher model by training a two-layer MLP to connect InterViT [14] and the Qwen2.5 LLM [77]. This teacher model is then used to perform understanding distillation for the Und. FFN and the Patch Embedding layer. Simultaneously, we perform generation pretraining to optimize the Gen. FFN. All other parameters of the LLM remain frozen to preserve its pretrained language capabilities. In Stage 2 and 3, the entire model is unfrozen to conduct unified mid-training and supervised fine-tuning (SFT), respectively. The VAE component for visual generation is omitted from the figure for clarity.
4.1 Stage-1: Multilmodal Pretraining
The objective of this stage is to equip the OneCAT with foundational visual perception and generation abilities while maintaining the linguistic capabilities from the pretrained LLM. The core challenge is that, for visual perception, the Und. FFN is initialized from the weights of the LLM’s text-focused FFN. While this "warm start" benefits abstract reasoning, it inherently lacks pretrained visual knowledge, making the training process highly data-intensive. To address this limitation, we leverage the visual perception capabilities of an MLLM teacher and introduce an understanding distillation strategy to optimize Und. FFN that significantly enhances visual learning efficiency. In parallel, we conduct generation pretraining to optimize Gen. FFN.
4.1.1 Stage 1-1: Teacher Training
Rather than employing an off-the-shelf MLLM as teacher (e.g., Qwen2.5-VL [3]), we construct a custom teacher model to ensure parameter consistency between the LLM backbones of the teacher and student models, thereby improving distillation efficiency. Specifically, the teacher is built by connecting a pre-trained vision encoder (InterViT [14]) and a LLM (Qwen2.5 [77]) with a two-layer MLP as connector. During this sub-stage, we freeze the ViT and LLM, and exclusively train the MLP connector on a curated small-scale dataset of image-to-text caption pairs using the standard NTP loss.
4.1.2 Stage 1-2: Expert Pretraining
With the teacher model prepared, we proceed to train the OneCAT model. We keep the QKV Projection, Attention, and Text FFN frozen, and selectively optimize the task-specific modules: the Und. FFN and Patch Embedding Layer for multimodal understanding, and the Gen. FFN for text-to-image generation.
Hyperparameter / Config Stage 1-1 Stage 1-2 Stage 2 Stage 3
(Teacher Training) (Expert Pretraining) (Unified Mid-Training) (Unified SFT)
Learning Rate $2\times 10^{-3}$ $2\times 10^{-4}$ $2\times 10^{-5}$ $1\times 10^{-5}$
LR Scheduler Cosine Cosine Cosine Cosine
Weight Decay 0 0 0.01 0.01
Gradient Norm Clip 1.0 1.0 1.0 1.0
Batch Size 512 2048 512 256
Sequence Length 1024 1024 8192 16384
Number of Sample: Text-Only - - 40M 2M
Number of Sample: Und. 10M 436M 70M 11M
Number of Sample: Gen. - 52M 60M 3M
Number of Token (Total) 5B 0.3T 0.6T 57B
Token Ratio (T:U:G): 0:1:0 0:8:1 1:2:6 1:5:6
Resolution: Und. $\times448$ $\times448$ Native Native
Use thumbnail $\times$ $\times$ $\checkmark$ $\checkmark$
Resolution: Gen. - $\times256$
Dynamical
[…]
Dynamical
[…]
Number of Scales : Gen. - 7 $10\sim$13
Table 1: Detailed hyperparameter and configuration of the training recipe across different stages.
Understanding Distillation: We optimize the Und. FFN on a large-scale dataset of image-to-text caption pairs to acquire fundamental visual knowledge. Based on the specifically designed teacher, the training objective is a combination of the NTP loss ($\mathcal{L}_{\text{NTP}}$) and a distillation loss ($\mathcal{L}_{\text{Distill}}$). Specifically, $\mathcal{L}_{\text{NTP}}$ is the standard cross-entropy loss for autoregressive text generation. For distillation, instead of matching the final output logits, we align the student’s internal hidden states with those of the teacher model through deep feature-level matching. This strategy enables the student to not only mimic the teacher’s final prediction but also its intermediate computational patterns across all token positions (including both visual and text tokens) for better visual knowledge transfer. The distillation loss $\mathcal{L}_{\text{Distill}}$ is formulated as the sum of MSE losses between the hidden states of the student and teacher models over all $N$ transformer layers:
$\mathcal{L}_{\text{Distill}}=\sum_{n=1}^{N}\text{MSE}(\mathbf{h}_{S}^{(n)},\mathbf{h}_{T}^{(n)}),$ (1)
where $\mathbf{h}_{S}^{(n)}$ and $\mathbf{h}_{T}^{(n)}$ represent the hidden state outputs from the $n$-th transformer block of the student and teacher models, respectively. The final objective is thus:
$\mathcal{L}_{\text{Und}}=\mathcal{L}_{\text{NTP}}+\lambda\mathcal{L}_{\text{Distill}},$ (2)
where $\lambda=0.2$ is a balancing hyperparameter. Throughout distillation, all models process images at a fixed resolution of $448\times 448$ to balance computational load and the granularity of visual features.
Generation Pretraining: In parallel, we optimize the Gen. FFN on a delicate text-to-image generation dataset to enable the model to learn the spatial relationships and dependencies between multi-scale discrete visual tokens. We adopt the cross-entropy loss for next-scale prediction [63, 24] and the output image resolution is fixed to $\times256$ in this stage.
4.2 Stage-2: Unified Mid-Training
In the second stage, we unfreeze the entire model to achieve unified mid-training across multiple tasks (i.e., image-to-text, text-to-image, image editing, and text-only dialogues). At this stage, we incorporate the proposed scale-aware adapter (SAA), which is optimized with other modules of LLM together to extract scale-specific representation for enhanced image generation.
We introduce native resolution strategy for both understanding and generation in this stage. For multimodal understanding, the model is trained to process images at their original resolutions, thereby preserving fine-grained details and eliminating information loss. Additionally, a thumbnail of resolution $\times448$ is included to provide global visual context. For visual generation, the model is trained to generate images with different resolution and aspect ratios where the side lengths dynamically sampled from a range of 288 to 864 pixels, which significantly enhancing its generative versatility and real-world applicability.
4.3 Stage-3: Unified Supervised Fine-tuning
The final stage involves unified supervised fine-tuning (SFT) using a curated dateset of higher-quality data to enhance instruction-following and visual generation quality. The native resolution strategy was continued, with the size of generated image expanded to support side lengths between 288 to 1728 pixels, enabling higher-resolution results.
5 Data Setup
Stage-1: For the multimodal understanding, we curate a large-scale dataset of approximately 436 million image-text pairs, which is meticulously compiled and processed through comprehensive filtering and deduplication. This dataset is collected from two primary sources: (1) Public Available Image-Text Caption Pairs: We incorporate several publicly available, high-quality image-caption datasets, including Recap-DataComp-1B [38], Capsfusion [81], Detailed-Caption [39], SA1B-Dense-Caption [15], and Moondream2-COYO-5M-Captions [31]. (2) Re-captioned Image Datasets: We generate new image-text pairs by re-captioning large-scale public image collections. The source image datasets for this process include COYO700M [6], CC12M [7], CC3M [58], LAION-400M [57], and Zeor250M [74]. From this dataset, we randomly select a subset of 10 million samples to train the custom teacher.
For image generation, we construct a dataset of 52 million text-to-image samples after a rigorous filtering process to remove samples with low resolution or poor aesthetic scores. This collection consists of 1 million class-labeled images from ImageNet-1k [17], 20 million pairs from re-captioned public collections (i.e., COYO700M [6], LAION-400M [57] and CC12M [7]), and 30 million in-house synthetic images generated by FLUX. The overall training token ratio across multimodal understanding and visual generation samples in Stage-I is approximately 8:1.
Stage-2: In the unified mid-training, for multimodal understanding we leverage an in-house dataset of 70 million visual instruction samples. This dataset is specifically curated to be highly diverse tasks, including general VQA, detailed image captioning, OCR, multimodal reasoning(i.e., STEM problem-solving), knowledge, and visual grounding. For visual generation, we supplement the text-to-image samples of Stage-1 with a additional collection of 8 million image editing samples, resulting a total of 60 million visual generation samples. These additional image editing samples are sourced from several public image editing datasets, including AnyEdit [80], UltraEdit [88], HQ-Edit [30] and OmniEdit [68]. Additionally, we incorporate 40 million text-only instruction samples to preserve the language ability of LLM. To ensure a strong focus on visual generation in Stage-II, we oversample the visual generation data, resulting a final training token ratio of approximately 1 :2 :6 across text-only, multimodal understanding, and visual generation tasks, respectively.
OneCAT-1.5B OneCAT-3B
Base Model Qwen2.5-1.5B-instruct Qwen2.5-3B-instruct
Active Parameters 1.5B 3B
Total Parameters 4.5B 9B
Understanding Distillation
Teacher ViT InterViT-300M
Teacher LLM Qwen2.5-1.5B-instruct Qwen2.5-3B-instruct
Table 2: Model configurations for the two variants of OneCAT.
Stage-3: In the SFT stage, for multimodal understanding and text-only instruction, we construct a high-quality dataset of 13 million samples, combining 10 million from public MAmmoTH-VL dataset [22] with 3 million of our in-house synthetic instruction samples to further improve the reasoning abilities. For visual generation, we utilize a total of 3 million samples, aggregated from UniWorld [44], blip3o-60k [8], ShareGPT-4o-Image [10], and additional in-house synthetic data generated by GPT-4o and FLUX. The overall training token ratio across text-only, multimodal understanding, and visual generation is approximately 1 :5 :6.
6 Implementation details
Model Configurations: We conduct our experiments on two model variants, OneCAT-1.5B and OneCAT-3B. The OneCAT-1.5B model is based on Qwen2.5-1.5B-instruct [3] and contains 1.5B active parameters (4.5B total). The OneCAT-3B model is built upon Qwen2.5-3B-instruct [3] and utilizes 3B active parameters (9B total). For understanding distillation, the ViT of teacher model is InterViT-300M [14] for both two variants and the LLM of teacher model is Qwen2.5-1.5B-instruct and Qwen2.5-3B-instruct for OneCAT-1.5B and OneCAT-3B, respectively, to align with the LLM backbone of our OneCAT. Detailed instruction of the model configurations is shown in Tab. 2.
Table 3: Performance comparison across multiple multimodal understanding benchmarks. Higher scores are better, as indicated by the up-arrow ($\uparrow$). A-LLM denotes the number of activated LLM parameters, while Vis. indicates the parameter count of the vision encoder or tokenizer for multimodal understanding. Chameleon [62] does not report the parameter count of its vision tokenizer. slash (/) denotes that models do not require a vision encoder or tokenizer for multimodal understanding. Top-1 accuracy is reported (Best in bold, second best is underlined).
[…]
A-LLM Vis.
Encoder-based Understanding Only Models
InternVL2-2B [13] 1.8B 0.3B 73.4 76.2 58.9 86.9 - 74.1
InternVL2.5-2B [12] 1.8B 0.3B 74.3 79.2 60.9 88.7 - 74.9
Qwen2-VL-3B [3] 1.5B 0.6B 79.7 73.5 65.5 90.1 - 74.7
Qwen2.5-VL-3B [3] 3B 0.6B 79.3 84.0 77.1 93.9 - 81.6
Encoder-free Understanding Only Models
Mono-InternVL [49] 1.8B / 72.6 73.7 - - 59.5 68.6
EvE [18] 7B / 56.8 59.1 - - 62.6 61.0
EvEv2 [19] 7B / 71.1 73.9 - - 62.9 74.8
HoVLE [61] 2.6B / 70.9 78.6 55.7 86.1 64.9 73.0
Unified Models
Chameleon [62] 7B - 4.8 2.9 5.0 1.5 - 46.0
Emu3 [67] 8B 0.3B 64.7 68.6 43.8 76.3 60.3 70.0
Harmon-1.5B [71] 1.5B 0.9B - - - - 58.9 -
Show-o2-1.5B [76] 1.5B 0.5B - - - - 60.0 69.0
Janus-Pro-1.5B [11] 1.5B 0.3B - - - - 59.3 -
[…]
ILLUME+ [28] 3B 0.6B - 69.9 44.1 80.8 - 74.2
VILA-U [72] 7B 0.4B 60.8 - - - 60.8 -
Janus-Pro-7B [11] 7B 0.3B - - - - 62.0 -
Tar-7B [23] 7B 0.4B - - - - 61.3 -
Show-o2-7B [76] 7B 0.5B - - - - 63.1 78.6
[…]
Data Packing and Gradient Accumulation: To optimize workload balance across distributed processes and increase training throughput, we employ a data packing strategy that concatenates multiple variable-length samples into contiguous sequences. Furthermore, to manage the gradient contributions and token ratios between modalities as in Table 1, we utilize an uneven gradient accumulation strategy: prior to each optimizer step, we accumulate a distinct number of micro-batches’ gradients for the text and image generation tasks to obtain a gradient of desired token ratios. Such an approach provides fine-grained control over the effective batch sizes of different tasks, ensuring a balanced and stable joint-training.
Unbiased Global Batch Gradients: When training on $N$ distributed processes, naively averaging local loss can lead to biased gradients when per-process token counts vary. The ideal objective is to optimize the Global Batch Loss, defined as the loss summed over tokens for all micro-batches, normalized by the global token count, denoted as $T_{global}$. To this end, we first prefetch all micro-batches for the next optimizer step, enabling each process to compute the local token counts; a subsequent All-Reduce collective operation then aggregates these local token counts into the final global token count, i.e., $T_{global}$. Similar to [42], we then employ Global Batch Reduced Loss by dividing each micro-batch loss by the averaged token count per process, $\tfrac{T_{global}}{N}$, which can be shown that the final synchronized gradient for the subsequent optimizer step is mathematically equivalent to the gradient of the global batch loss, enabling training with unbiased gradients.
Table 4: Performance comparison across multiple multimodal understanding benchmarks. Higher scores are better, as indicated by the up-arrow ($\uparrow$). A-LLM denotes the number of activated LLM parameters, while Vis. indicates the parameter count of the vision encoder or tokenizer for multimodal understanding. Chameleon [62] does not report the parameter count of its vision tokenizer. slash (/) denotes that models do not require a vision encoder or tokenizer for multimodal understanding. Top-1 accuracy is reported (Best in bold, second best is underlined).
[…]
A-LLM Vis.
Encoder-based Understanding Only Models
InternVL2 [13] 1.8B 0.3B 1440 1877 73.2 34.3 44.6 46.4 71.6
InternVL2.5 [12] 1.8B 0.3B - 2138 74.7 43.6 60.8 51.3 -
Qwen2-VL [65] 1.5B 0.6B - 1872 74.9 41.1 49.5 43.0 -
Qwen2.5-VL [3] 3B 0.6B - 2157 79.1 53.1 61.8 62.3 -
Encoder-free Understanding Only Models
Mono-InternVL [49] 1.8B / - 1875 65.5 33.7 40.1 45.7 67.4
EvE [18] 7B / - 1628 52.3 32.6 25.7 - 64.6
EvEv2.0 [19] 7B / - 1709 66.3 39.3 45.0 - 71.4
HoVLE [61] 2.6B / - 1864 71.9 33.7 44.3 46.2 70.7
Unified Models
Chameleon [62] 7B - - - 35.7 28.4 8.3 - 30.6
Emu3 [67] 8B 0.3B - - 58.5 31.6 37.2 - 68.2
Harmon [71] 1.5B 0.9B 1155 1476 65.5 38.9 - - 67.1
Show-o2 [76] 1.5B 0.5B 1450 - 67.4 37.1 - - 65.6
Janus-Pro [11] 1.5B 0.3B 1444 - 75.5 36.3 39.8 - -
[…]
ILLUME+ [28] 3B 0.6B $14$ - 80.8 44.3 40.3 - 73.3
VILA-U [72] 7B 0.4B 1401 - - - 33.5 - 59.0
Janus-Pro [11] 7B 0.3B 1567 - 79.2 41.0 50.0 - -
Tar [23] 7B 0.4B 1571 1926 74.4 39.0 - - -
Show-o2 [76] 7B 0.5B 1620 - 79.3 48.9 - - 69.8
[…]
Figure 6: Text-to-Image comparison.
Figure 7: Image-Editing comparison.
7 Evaluation
7.1 Multimodal understanding.
We evaluate our model on 12 public Multimodal understanding benchmarks spanning diverse capabilities: MMbench [47], MME [79], MMMU [83], MM-Vet [82], and SEED [35] assess general multimodal perception and reasoning. MathVista [48] focuses on mathematical reasoning. TextVQA [60], ChartQA [50], InfoVQA [52], and DocVQA [51] evaluate OCR and text-related visual question answering; while GQA [29] evaluates visual scene understanding and AI2D [33] evaluates scientific diagram comprehension.
Table 5: Performance comparison on the GenEval [21] benchmark. The dagger () indicates methods that employ an LLM for prompt rewriting. Best in bold, second best is underlined.
[…]
Generation-only Models
SDXL [54] 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E 3 [59] 0.96 0.87 0.47 0.83 0.43 0.45 0.67
Infinity [24] - 0.85 - - 0.49 0.57 0.73
SD3-Medium [20] 0.99 0.94 0.72 0.89 0.33 0.60 0.74
FLUX.1-dev [34] 0.98 0.93 0.75 0.93 0.68 0.65 0.82
Unified Models
Chameleon-7B [62] - - - - - - 0.39
Transfusion-7B [90] - - - - - - 0.63
Emu3-8B [67] 0.99 0.81 0.42 0.80 0.49 0.45 0.66
ILLUME+ 3B [28] 0.99 0.88 0.62 0.84 0.42 0.53 0.72
Harmon-1.5B [71] 0.99 0.86 0.66 0.85 0.74 0.48 0.76
Show-o2-7B [76] 1.00 0.87 0.58 0.92 0.52 0.62 0.76
Janus-Pro-7B [11] 0.99 0.89 0.59 0.90 0.79 0.66 0.80
Mogao-7B [41] 1.00 0.97 0.83 0.93 0.84 0.80 0.89
BLIP3-o-8B [8] - - - - - - 0.84
Tar-7B [23] 0.99 0.92 0.83 0.85 0.80 0.65 0.84
UniWorld-V1-20B [43] 0.99 0.93 0.79 0.89 0.49 0.70 0.80
UniWorld-V1-20B [43] 0.98 0.93 0.81 0.89 0.74 0.71 0.84
BAGEL-7B [16] 0.99 0.94 0.81 0.88 0.64 0.63 0.82
BAGEL-7B [16] 0.98 0.95 0.84 0.95 0.78 0.77 0.88
[…]
As shown in Tab. 3 and Tab. 4, we compare OneCAT with three types of models: encoder-based understanding-only models, encoder-free understanding-only models, and unified MLLMs. Our OneCAT-3B model demonstrates superior performance, significantly outperforming all existing encoder-free understanding-only MLLMs, e.g., HoVLE [61] and EvEv2 [19], across nearly all benchmarks. For instance, on OCR-related tasks including TextVQA (73.9), ChartQA (81.2), InfoVQA (64.8), and DocVQA (91.2), OneCAT-3B achieves new state-of-the-art results among encoder-free models. It also excels in general vision-language benchmarks such as MME-P (1630), MMBench (78.8), and MM-Vet (52.2).
Moreover, OneCAT-3B outperforms recent unified MLLMs that rely on external vision encoders or tokenizers—such as Janus-Pro-7B [11] (using SigLIP [84]) and Tar-7B [23] (using SigLip2 [64])—despite activating fewer parameters. Compared to top-tier encoder-based understanding-only models like Qwen2.5-VL-3B [3], our model exhibits a slight performance gap, which we primarily attribute to differences in the scale and quality of training data. Specifically, Qwen2.5-VL was trained on 4T tokens, whereas our OneCAT was trained on only 0.5T tokens for multimodal understanding. We believe this gap can be bridged in the future by scaling up the pretraining data and incorporating more higher-quality instruction data.
7.2 Visual Generation.
We evaluate our model on three widely-used visual generation benchmarks: two for text-to-image generation, GenEval [21] and DPG-Bench [27], and one for instruction-based image editing, ImgEdit [78]. We follow previous works [10, 41, 16] to use Classifier-free guidance(CFG) [25] to enhance visual generation quality. During training, we randomly drop tokens of conditional text and reference image with probabilities 0.1, 0.1, respectively. During inference, we combines conditional and unconditional predicted logits to produce outputs that better adhere to the given conditions. To ensure a fair comparison, we adhere strictly to the original raw prompts for the GenEval benchmark, unlike some previous works that employ LLM-based prompt rewriting to enhance performance. As shown in Tab.5, 6, and 7, our OneCAT-3B model demonstrates highly competitive performance across all tasks.
On GenEval (Tab. 5), which evaluates fine-grained instruction following over object counts, colors, and spatial relationships, OneCAT-3B achieves a SOTA overall score of 0.90. This performance surpasses most all unified models, including the strong baseline BAGEL-7B (0.88 with prompt rewriting) and Mogao-7B (0.89 with prompt rewriting). Notably, OneCAT-3B excels in challenging categories such as Position and Color Attribute, where it achieves the best performance (0.84 and 0.80 respectively), showcasing its superior ability to interpret complex spatial and attribute-based instructions.
On DPG-Bench (Tab. 6), a benchmark focused on compositional text-to-image generation, OneCAT-3B attains a strong overall score of 84.53. This result is highly competitive among unified models, outperforming strong counterparts like Janus-Pro-7B (84.19) and Mogao-7B (84.33).
On ImgEdit-Bench (Tab. 7), a challenging and diverse image editing benchmark, OneCAT-3B achieves an overall score of 3.43. This places it firmly among the top-performing unified models and significantly outperforms many specialized editing models. OneCAT-3B demonstrates exceptional capabilities in categories requiring precise local and global adjustments, securing the top scores in Adjust (3.70), Extract (2.42), and Background (3.79) manipulation. This highlights the effectiveness of our model’s ability to condition its generation on fine-grained visual cues from a reference image.
Figures 6 and 7 present qualitative comparisons for the text-to-image and image-editing tasks, respectively. OneCAT-3B exhibits leading instruction-following and world-understanding capabilities among open-source models. For example, in the fourth row of Figure 6, it is the only open-source model that correctly generates exactly four characters in a pixel-art style; in contrast, BAGEL-7B produces five characters, and Janus-Pro-7B fails to render pixel art. Similarly, in the last row of Figure 7, only OneCAT-3B produces an image with the correct lighting conditions.
In summary, the strong instruction-following capability in both text-to-image generation and image editing tasks highlights the effectiveness and versatility of our architectural design.
Table 6: Performance comparison on the DPG-Bench [27] benchmark. Best in bold, second best is underlined.
[…]
Generation-only Models
Hunyuan-DiT [40] 84.59 80.59 88.01 74.36 86.41 78.87
Playground v2.5 [36] 83.06 82.59 81.20 84.08 83.50 75.47
PixArt-$\Sigma$ [9] 86.89 82.89 88.94 86.59 87.68 80.54
DALL-E 3 [59] 90.97 89.61 88.39 90.58 89.83 83.50
Infinity [24] 93.11 - - 90.76 - 83.46
SD3-Medium [20] 87.90 91.01 88.83 80.70 88.68 84.08
FLUX.1-dev [34] 82.10 89.50 88.80 91.10 89.40 84.00
Unified Models
Emu3-8B [67] - - - - - 81.60
Janus-Pro-7B [11] 86.90 88.90 89.40 89.32 89.48 84.19
Mogao-7B [41] 82.37 90.03 88.26 93.18 85.40 84.33
BLIP3-o-8B [8] - - - - - 81.60
Tar-7B [23] 83.98 88.62 88.05 93.98 84.86 84.19
Show-o2-7B [76] 89.00 91.78 89.96 91.81 91.64 86.14
[…]
Table 7: Comprehensive comparison on ImgEdit-Bench [78] showing performance across nine editing categories. Higher scores are better for all metrics. Best in bold, second best is underlined.
Model Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
Editing-only Models
MagicBrush [85] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.22 1.90
Instruct-Pix2Pix [5] 2.45 1.83 1.44 2.01 1.50 1.44 3.55 1.20 1.46 1.88
AnyEdit [32] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.65 2.45
UltraEdit [87] 3.44 2.81 2.13 2.96 1.45 2.83 3.76 1.91 2.98 2.70
Step1X-Edit [46] 3.88 3.14 1.76 3.40 2.41 3.16 4.63 2.64 2.52 3.06
ICEdit [86] 3.58 3.39 1.73 3.15 2.93 3.08 3.84 2.04 3.68 3.05
Unified Models
OmniGen [73] 3.47 3.04 1.71 2.94 2.43 3.21 4.19 2.24 3.38 2.96
OmniGen2 [70] 3.57 3.06 1.77 3.74 3.20 3.57 4.81 2.52 4.68 3.44
BAGEL-7B [16] 3.56 3.31 1.70 3.30 2.62 3.24 4.49 2.38 4.17 3.20
UniWorld-V1-20B [43] 3.82 3.64 2.27 3.47 3.24 2.99 4.21 2.96 2.74 3.26
OneCAT-3B 3.65 3.70 2.42 3.92 3.00 3.79 4.61 2.23 3.53 3.43
7.3 Comparison of Inference Efficiency
In addition to its strong performance, our model’s architectural design also yields significant improvements in inference efficiency for handling both high-resolution image input and output. We evaluated the efficiency of OneCAT-3B in two distinct phases: prefilling and generation.
In the prefilling phase, we measure the Time to the First Token (TTFT) to assess the computational cost of processing the input. As shown in Tab. 8, thanks to our pure decoder-only architecture that eliminates the need for a separate ViT encoder, OneCAT-3B demonstrates a substantial efficiency advantage, particularly when handling high-resolution images. Compared to QwenVL-3B, as the image resolution increases from $768\times 768$ to $1792\times 1792$, the reduction in TTFT for OneCAT-3B grows sharply from 50.4% to 61.4%. This provides strong evidence for the effectiveness of our design, especially for high-resolution inputs.
In the generation phase, we evaluated the total inference time for text-to-image (T2I) generation and image editing. As detailed in Tab. 9, OneCAT-3B’s inference speed far surpasses that of the strong diffusion-based baseline, BAGEL-7B. For instance, when generating a $1024\times 1024$ resolution image, OneCAT-3B’s T2I and editing inference times are merely 2.85s and 4.61s, respectively, making it approximately $10\times$ faster than BAGEL-7B. This significant advantage stems from our pioneering multi-scale autoregressive generation mechanism inside LLM that enables parallel token generation, which drastically reduces the number of required decoding steps compared to iterative diffusion models.
In summary, OneCAT achieves exceptional inference efficiency in both the prefilling and generation stages. This highlights the immense potential and practical value of our unified autoregressive framework for building efficient and powerful large multimodal models.
Table 8: Efficiency comparison of OneCAT-3B and QwenVL-3B. Models are tested based on one NVIDIA H800 GPU. We report the time to the first token (TTFT) to measure the reduction of computational cost of prefilling phase. 256* denotes the number of visual tokens for thumbnail.
Model Resolution of Input Image #Input Text Tokens #Input Visual Tokens TTFT(s) Reduction
Qwen2.5-VL-3B $768\times 768$ 24 731 0.135 50.4%
OneCAT-3B $768\times 768$ 24 731 $256^{*}$ 0.067
Qwen2.5-VL-3B $1024\times 1024$ 24 1395 0.216 57.4%
OneCAT-3B $1024\times 1024$ 24 1395 $256^{*}$ 0.092
Qwen2.5-VL-3B $1792\times 1792$ 24 4098 0.583 61.4%
OneCAT-3B $1792\times 1792$ 24 4098 $256^{*}$ 0.225
Table 9: Generation efficiency comparison of OneCAT-3B and BAGEL. Models are tested based on one NVIDIA H800 GPU. We report total inference time to measure the Text-to-Image(T2I) and Image-Editing efficiency.
Model Resolution of Generated Image T2I Infer. Time (s) Edit Infer. Time (s)
BAGEL-7B $512\times 512$ 8.76 13.45
OneCAT-3B $512\times 512$ 1.40 2.03
BAGEL-7B $1024\times 1024$ 26.29 46.44
OneCAT-3B $1024\times 1024$ 2.85 4.61
8 Conclusion
In this work, we presented OneCAT, a pure decoder-only unified multimodal model that seamlessly integrates understanding, generation, and editing within a single, streamlined architecture. By eliminating external encoders and tokenizers, employing a modality-specific MoE design, and introducing a multi-scale autoregressive generation mechanism, OneCAT achieves strong performance across a wide range of benchmarks while significantly improving inference efficiency. Our results demonstrate the viability and advantages of a first-principles approach to multimodal modeling, offering a powerful new baseline for future research and applications in general-purpose multimodal intelligence.

### v2

@@
 We contend that a true paradigm shift—necessary to unlock the full potential of unified systems—demands a move towards a more fundamental, first-principles architecture that eschews external components.
-We therefore propose that a pure decoder-only transformer, trained under a unified objective, provides not just a sufficient, but a more elegant and potent foundation for the next generation of general-purpose multimodal intelligence.
+We therefore propose a pure decoder-only transformer, trained under a unified objective, providing not just a sufficient, but a more elegant and potent foundation for the next generation of general-purpose multimodal intelligence.
 Figure 1: Showcase of the text-to-image generation abilities of the OneCAT model.
@@
 For example, Fuyu-8B [4] processes vision patches by feeding them through a simple linear patch embedding layer directly into the LLM backbone, which markedly reduce inference latency.
-Inspired by this success, subsequent works [18, 49, 19] further advance decode-only MLLMs by targeting their training processes and architectures.
-EvE [18] introduces a novel training objective that guides the model’s visual learning by aligning the last layer of LLM’s hidden states of image patch tokens with semantic features from a pre-trained vision encoder.
+Inspired by this success, subsequent works [18, 51, 19, 35, 68, 62], further advance decode-only MLLMs by targeting their training processes and architectures.
+EvE [18] and VoRA [68] aligns the LLM’s hidden states with semantic features from a pre-trained ViT.
+However, directly using a smaller model (e.g., a ViT with hundreds of millions of parameters) as the teacher to distill knowledge into a significantly larger LLM (with several billion parameters) may restrict the expressive capacity of the LLM.
 Differently, Mono-InternVL [51] and EvEv2.0 [19] adopt a Mixture-of-Experts (MoE) framework, introducing a dedicated visual expert to handle visual-specific features more effectively.
@@
 2.3 Unified Visual Understanding and Generation
-Building on the success of MLLMs, the convergence of visual understanding and generation into a unified framework now represents a key research frontier.
-Pioneering unified MLLMs such as Chameleon [62],Transfusion [90], emu3 [67] and show-o [75] utilize visual tokenizers (e.g., VQ-VAE) to convert images into a sequence of discrete tokens, enabling seamless multimodal understanding and generation within a single model.
+Building on the success of MLLMs, the convergence of visual understanding and generation into a unified framework now represents a key research frontier [73, 11, 79, 95, 71, 80, 38, 55, 8, 45, 16, 43].
+Pioneering unified MLLMs such as Chameleon [65],Transfusion [95], emu3 [71], show-o [80] and Synergen-vl [38] utilize visual tokenizers (e.g., VQ-VAE) to convert images into a sequence of discrete tokens, enabling seamless multimodal understanding and generation within a single model.
 However, the discretization inevitably results in lossy visual information and weakens in extracting semantic contents.
@@
 Figure 5: Overview of the training pipeline.
-In Stage 1, we first prepare a teacher model by training a two-layer MLP to connect InterViT [14] and the Qwen2.5 LLM [77].
+In Stage 1, we first prepare a teacher model by training a two-layer MLP to connect InternViT [14] and the Qwen2.5 LLM [82].
 This teacher model is then used to perform understanding distillation for the Und.
@@
 Rather than employing an off-the-shelf MLLM as teacher (e.g., Qwen2.5-VL [3]), we construct a custom teacher model to ensure parameter consistency between the LLM backbones of the teacher and student models, thereby improving distillation efficiency.
-Specifically, the teacher is built by connecting a pre-trained vision encoder (InterViT [14]) and a LLM (Qwen2.5 [77]) with a two-layer MLP as connector.
+Specifically, the teacher is built by connecting a pre-trained vision encoder (InternViT [14]) and a LLM (Qwen2.5 [82]) with a two-layer MLP as connector.
 During this sub-stage, we freeze the ViT and LLM, and exclusively train the MLP connector on a curated small-scale dataset of image-to-text caption pairs using the standard NTP loss.
@@
 Understanding Distillation
-Teacher ViT InterViT-300M
+Teacher ViT InternViT-300M
 Teacher LLM Qwen2.5-1.5B-instruct Qwen2.5-3B-instruct
@@
 The OneCAT-3B model is built upon Qwen2.5-3B-instruct [3] and utilizes 3B active parameters (9B total).
-For understanding distillation, the ViT of teacher model is InterViT-300M [14] for both two variants and the LLM of teacher model is Qwen2.5-1.5B-instruct and Qwen2.5-3B-instruct for OneCAT-1.5B and OneCAT-3B, respectively, to align with the LLM backbone of our OneCAT.
+For understanding distillation, the ViT of teacher model is InternViT-300M [14] for both two variants and the LLM of teacher model is Qwen2.5-1.5B-instruct and Qwen2.5-3B-instruct for OneCAT-1.5B and OneCAT-3B, respectively, to align with the LLM backbone of our OneCAT.
 Detailed instruction of the model configurations is shown in Tab.
@@
 slash (/) denotes that models do not require a vision encoder or tokenizer for multimodal understanding.
-Top-1 accuracy is reported (Best in bold, second best is underlined).
+Best in bold, second best is underlined (across unified models).
 […]
@@
 EvEv2 [19] 7B / 71.1 73.9 - - 62.9 74.8
+VoRA [68] 7B / 56.3 - - - - 65.6
+SAIL [35] 7B / 77.1 - - - - 76.7
 HoVLE [64] 2.6B / 70.9 78.6 55.7 86.1 64.9 73.0
@@
 EvEv2.0 [19] 7B / - 1709 66.3 39.3 45.0 - 71.4
+VoRA [68] 7B / 1363 1674 64.2 32.2 33.7 - 64.2
+SAIL [35] 7B / - 1719 70.1 - 46.3 57.0 72.9
 HoVLE [64] 2.6B / - 1864 71.9 33.7 44.3 46.2 70.7
@@
 Our OneCAT-3B model demonstrates superior performance, significantly outperforming all existing encoder-free understanding-only MLLMs, e.g., HoVLE [64] and EvEv2 [19], across nearly all benchmarks.
-For instance, on OCR-related tasks including TextVQA (73.9), ChartQA (81.2), InfoVQA (64.8), and DocVQA (91.2), OneCAT-3B achieves new state-of-the-art results among encoder-free models.
+For instance, on OCR-related tasks including AI2D (77.8), ChartQA (81.2), InfoVQA (64.8), and DocVQA (91.2), OneCAT-3B achieves new state-of-the-art results among encoder-free models.
 It also excels in general vision-language benchmarks such as MME-P (1630), MMBench (78.8), and MM-Vet (52.2).
@@
 This highlights the immense potential and practical value of our unified autoregressive framework for building efficient and powerful large multimodal models.
-Table 8: Efficiency comparison of OneCAT-3B and QwenVL-3B.
+Table 8: Efficiency comparison of OneCAT-3B and Qwen2.5-VL-3B.
 Models are tested based on one NVIDIA H800 GPU.

### v3

@@
 Abstract
-We introduce OneCAT, a unified multimodal model that seamlessly integrates understanding, generation, and editing within a novel, pure decoder-only transformer architecture.
-Our framework uniquely eliminates the need for external components such as Vision Transformers (ViT) or vision tokenizer during inference, leading to significant efficiency gains, especially for high-resolution inputs.
-This is achieved through a modality-specific Mixture-of-Experts (MoE) structure trained with a single autoregressive (AR) objective, which also natively supports dynamic resolutions.
-Furthermore, we pioneer a multi-scale visual autoregressive mechanism within the Large Language Model (LLM) that drastically reduces decoding steps compared to diffusion-based methods while maintaining state-of-the-art performance.
+We introduce OneCAT, a unified multimodal model that seamlessly integrates understanding, generation, and editing within a single decoder-only transformer architecture.
+OneCAT uniquely eliminates the need for external components such as Vision Transformers (ViT) or vision tokenizer during inference, leading to significant efficiency gains, especially for high-resolution image inputs and outputs.
+This is achieved through a modality-specific Mixture-of-Experts (MoE) design trained with a unified autoregressive (AR) objective, which also natively supports dynamic resolutions.
+Furthermore, we pioneer to achieve multi-scale visual autoregressive mechanism within the Large Language Model (LLM) with proposed scale-aware adapter (SAA) that drastically reduces decoding latency compared to diffusion-based methods while maintaining state-of-the-art performance.
 Our findings demonstrate the powerful potential of pure autoregressive modeling as a sufficient and elegant foundation for unified multimodal intelligence.
-As a result, OneCAT sets a new performance standard, outperforming existing open-source unified multimodal models across benchmarks for multimodal generation, editing, and understanding.
+As a result, OneCAT outperforms existing unified models across benchmarks for multimodal understanding, generation, and editing.
+[…]
+[…]
 1 Introduction
-In the past few years, modular approaches, utilizing separate architectures for understanding [2, 69, 3, 13, 14, 12], generation [42, 37, 61, 24, 34, 20], and editing tasks [90, 5, 32, 92, 48, 91], dominated multi-modal frameworks.
-While producing capable systems, it inherently creates complex, multi-stage pipelines.
-Such designs often suffer from architectural bottlenecks that limit deep, early-stage fusion of cross-modal information and introduce significant inference latency, presenting a major barrier to both efficiency and performance.
-In response to these limitations, the field has seen a rapid convergence towards unified multimodal modeling, aiming to integrate these disparate capabilities within a single, end-to-end architecture [71, 65, 80, 81, 11, 16].
-Despite the trend towards unification, many current models remain tethered to the modular paradigm.
-We contend that a true paradigm shift—necessary to unlock the full potential of unified systems—demands a move towards a more fundamental, first-principles architecture that eschews external components.
-We therefore propose a pure decoder-only transformer, trained under a unified objective, providing not just a sufficient, but a more elegant and potent foundation for the next generation of general-purpose multimodal intelligence.
+Modular approaches using separate modules for understanding [92, 5, 17], generation [81, 45, 25], and editing tasks [114, 7, 115] are becoming dominant for multi-modal frameworks.
+Despite producing capable systems, such designs inherently create complex multi-stage pipelines suffering from architectural bottlenecks that limit deep, early-stage fusion of cross-modal information and introduce significant inference latency, presenting a major barrier to both efficiency and performance.
+Unified multimodal LLMs aim to address these limitations by integrating these disparate abilities within a single end-to-end architecture [94, 15, 21], but many methods remain tethered to the modular paradigm [73, 11].
+This motivates us to achieve a true revolution on the fundamental architecture that unlocks the full potential of unified systems and eschews heavy external components.
+In this paper, we propose a single decoder-only autoregressive model trained under a unified objective to provide elegant and potent foundation for general-purpose multimodal intelligence.
+Unified architecture.
+We propose the first encoder-free framework for unified multimodal LLM (MLLM), where raw visual inputs are directly tokenized into patch embeddings and are processed alongside text tokens within a single decoder model.
+The critical innovation is a modality-specific MoE layer that dynamically routes continuous vision tokens, discrete vision tokens, and text tokens to specialized experts.
+It enables deep and early-stage feature fusion without requiring exquisite encoders (i.e., ViT and vision tokenizer) for efficient inference.
+For generative tasks, we pioneer to embed the multi-scale autoregressive mechanism [89] into the LLM and propose a scale-aware adapter (SAA) to extract scale-specific representation for augmentation.
+Thus, image tokens can be predicted from low to high resolutions, while next-token prediction is adopted for text tokens.
+This design significantly enhances the speed and quality of image generation by simultaneously circumventing the high latency of diffusion models and learning a coarse-to-fine generative process.
 Figure 1: Showcase of the text-to-image generation abilities of the OneCAT model.
 Figure 2: Showcase of the image editing abilities of the OneCAT model, including general image editing tasks such as object removal, background adjustment, color adjustment, subject replacement, and style transfer; as well as perceptual tasks including depth estimation, pose estimation, object segmentation, and Canny edge detection.
-To realize this vision, we re-evaluate the core architectural tenets of multimodal systems.
-We introduce an encoder-free framework where raw visual inputs are directly tokenized into patch embeddings and processed alongside text tokens within a single decoder stack.
-The critical innovation is a modality-specific Mixture-of-Experts (MoE) layer, which dynamically routes vision and text tokens to specialized experts, enabling deep, early-stage feature fusion in an efficient inference manner without the need for exquisite encoders.
-For generative tasks, we pioneer a multi-scale autoregressive mechanism [24, 66] inside the LLM, augmented with the proposed scale-aware adapter modules, to predict image tokens from low to high resolution.
-This design not only circumvents the high latency of iterative diffusion models but also allows the model to learn a coarse-to-fine generative process, significantly enhancing both speed and output quality.
-Regarding data protocol, we leverage mixed training strategies that combine large-scale, web-scraped image-text pairs with curated, instruction-following datasets.
-Our pre-training corpus is constructed to expose the model to a diverse range of data formats, including image-text documents, visual question-answering pairs, and generative prompts for both text-to-image synthesis and image editing.
-This heterogeneous data mixture is crucial for training a truly unified model, as it forces the shared decoder to develop a generalized representation that can seamlessly switch between comprehension, generation, and editing tasks.
-We argue that this unified data diet, coupled with our bottleneck-free architecture, is key to fostering the synergistic emergence of complex multimodal capabilities.
-Building upon these principles, we present Only DeCoder Auto-regressive Transformer (OneCAT), an open-source unified multimodal model.
-Comprehensive evaluations demonstrate that OneCAT sets a new state-of-the-art for pure decoder-only unified models.
-More importantly, the encoder-free design provides a significant inference speedup, particularly for high-resolution inputs.
-By demonstrating the viability and superiority of a pure decoder architecture, OneCAT offers a more first-principles-aligned paradigm for multimodal modeling.
-It facilitates earlier cross-modal fusion through its unified MoE structure and enhances semantic consistency, providing valuable insights and a powerful new baseline for the development of next-generation unified multimodal systems.
+Unified training paradigm.
+Training encoder-free MLLMs for strong visual perception ability is notoriously data-intensive [64].
+To mitigate this, existing methods like VoRA [91] and EvE [23] perform distillation to align the internal hidden states of an LLM student with a pre-trained ViT teacher.
+However, they suffer from supervisory bottleneck problem that the expressive capacity of an LLM is restricted by a small teacher.
+We address this problem with a novel distillation strategy that first customizes a powerful MLLM teacher and then efficiently transfers its comprehensive visual perception to the proposed encoder-free unified MLLM.
+On such basis, we use large-scale heterogeneous multimodal data and employ a unified expert pretraining, mid-training, and supervised fine-tuning to force the shared decoder to achieve generalized representation that can seamlessly switch between comprehension, generation, and editing tasks.
+Building upon these innovations, we present Only DeCoder Auto-regressive Transformer (OneCAT), a unified multimodal model.
+Comprehensive evaluations demonstrate that OneCAT sets a new state-of-the-art for unified models.
+More importantly, the encoder-free design provides a significant inference speedup, particularly for high-resolution inputs and outputs.
+OneCAT demonstrates the viability and superiority of a pure decoder architecture, and offers a more first-principle-aligned paradigm for multimodal modeling.
+It facilitates earlier cross-modal fusion through its unified MoE structure and enhances semantic consistency via unified AR objective to provide valuable insights and a powerful new baseline for developments of next-generation unified multimodal systems.
 For examples of our model’s impressive image generation capabilities, please refer to Fig.
@@
 The field of Multimodal Large Language Models (MLLMs) has rapidly evolved, converging on a dominant compositional architecture.
-This paradigm connects a pre-trained vision encoder, such as CLIP [58], SigLIP [89], or the more recent InternViT [14], to a powerful Large Language Model (LLM) through a trainable connector.
+This paradigm connects a pre-trained vision encoder (i.e., CLIP [77], SigLIP [113], and InternViT [19]), to a powerful LLM through a trainable connector.
 Pioneering works [3, 51] propose sophisticated connector designs.
-For example, Flamingo [1] introduces gated cross-attention layers to inject visual information into an LLM, while BLIP-2 [39] develops the Q-Former to bridge the modality gap between an image encoder and an LLM.
-A significant shift occurs with LLaVA [47], which simplifies the connector to a lightweight Multi-Layer Perceptron (MLP) projection layer.
-This effective design become a foundational blueprint for numerous subsequent MLLMs.
-For example, recent state-of-the-art models like the InternVL series [14, 12, 13, 96, 70] and the Qwen-VL series [69, 3, 2] adopt this same core principle and achieve superior performance by leveraging larger-scale training data and more powerful vision and language foundation models.
+For example, Flamingo [3] introduces gated cross-attention layers to inject visual information into an LLM, while BLIP-2 [51] develops the Q-Former to bridge the modality gap between the vision encoder and LLM.
+A significant shift occurs with LLaVA [60], which simplifies the connector to a lightweight MLP, which become a foundational blueprint for subsequent MLLMs.
+For example, recent state-of-the-art models like the InternVL series [19, 17, 18, 120, 93] and the Qwen-VL series [92, 5, 4] adopt this core principle and achieve superior performance by leveraging larger-scale training data and more powerful vision and language foundation models.
 However, this successful compositional design has inherent drawbacks.
 The separate nature of the vision and language components complicates the end-to-end optimization process and introduces two critical bottlenecks.
-First, the sequential nature of the architecture, where the vision encoder must fully process an image before the LLM can begin its generation, leads to high inference latency.
-This initial step is often referred to as the “prefilling” stage.
-Second, the connector itself acts as an information bottleneck.
+First, the sequential nature of the architecture, where the vision encoder must fully process an image before the LLM can begin its generation, leads to high inference latency, especially for the prefilling stage.
+Second, the connector acts as an information bottleneck.
 In this so-called late fusion pipeline, complex visual information is compressed into a compact representation for the LLM, inevitably causing a loss of fine-grained visual detail.
@@
 2.2 Decoder-only MLLMs
-Decoder-only MLLMs, also known as monolithic MLLMs, have recently emerged as a minimalist yet powerful alternative to the conventional compositional architecture.
-This paradigm aims to achieve greater efficiency and a more direct early fusion of modalities by eliminating the separate vision encoder.
-For example, Fuyu-8B [4] processes vision patches by feeding them through a simple linear patch embedding layer directly into the LLM backbone, which markedly reduce inference latency.
+Decoder-only MLLMs, also known as monolithic MLLMs, have recently emerged as a minimalist yet powerful alternative to the compositional MLLMs.
+This paradigm aims to achieve greater efficiency and a more direct early fusion of modalities by removing the separate vision encoder or tokenizer.
+For example, Fuyu-8B [6] processes vision patches by feeding them through a simple linear patch embedding layer directly into the LLM, which markedly reduce inference latency.
 Inspired by this success, subsequent works [23, 64, 24, 47, 91, 82], further advance decode-only MLLMs by targeting their training processes and architectures.
@@
 HoLVE [87] prepends a causal transformer to the LLM to explicitly convert both visual and textual inputs into a shared space.
-Despite these promising advancements, the overall training efficiency of these models remains a significant challenge.
+Despite these promising advancements, the overall training efficiency of decoder-only MLLMs remains a significant challenge.
 More importantly, the potential for the decoder-only architecture to create unified models that can seamlessly integrate multimodal understanding, generation, and even image editing capabilities remains a largely unexplored research avenue.
@@
 Building on the success of MLLMs, the convergence of visual understanding and generation into a unified framework now represents a key research frontier [96, 15, 102, 119, 94, 103, 50, 73, 11, 57, 21, 55].
-Pioneering unified MLLMs such as Chameleon [65],Transfusion [95], emu3 [71], show-o [80] and Synergen-vl [38] utilize visual tokenizers (e.g., VQ-VAE) to convert images into a sequence of discrete tokens, enabling seamless multimodal understanding and generation within a single model.
+Pioneering unified MLLMs such as Chameleon [88],Transfusion [119], emu3 [94], show-o [103] and Synergen-VL [50] utilize vison tokenizer (e.g., VQ-VAE) to convert images into discrete tokens, enabling seamless multimodal understanding and generation within a single model.
 However, the discretization inevitably results in lossy visual information and weakens in extracting semantic contents.
@@
 However, since the visual tokens (i.e., the input of the visual decoder) must be decoded token-by-token through the LLM before subsequent next-scale prediction, this approach compromises the inference efficiency that is the key advantage of the NSP.
-In contrast, our proposed OneCAT seamlessly unifies next-token prediction for text generation and next-scale prediction for visual generation within a single LLM.
+In contrast, our proposed OneCAT seamlessly unifies next-token prediction for text generation and next-scale prediction for visual generation within a single decoder-only transformer of the LLM, and proposes the scale-aware adapter to further exploit the scale-specific representation for enhanced visual generation.
 3 OneCAT
-Figure 3: Overview inference pipeline of OneCAT.
-OneCAT is a decoder-only autoregressive unified model that seamlessly supports multimodal understanding, text-to-image generation and image editing.
-OneCAT implements both next-token prediction for text generation and next-scale prediction for visual generation within a unified LLM backbone.
-In the left part, the dash squares denote the generated text and visual token during inference.
-In the right part, PE and $T$ denote patch embedding and text tokenizer, respectively.
-TD and ID denote text detokenizer and image detokenizer, respectively.
-As illustrated in Fig.
-3, OneCAT employs a pure decoder-only architecture, eliminating the need for any additional vision encoder or tokenizer during inference.
-This streamlined design significantly simplifies the model structure and reduces computational overhead.
-Unlike traditional multimodal models that rely on semantic encoders like ViTs for understanding [3, 16, 43, 11], OneCAT utilizes a lightweight Patch Embedding layer.
-This layer losslessly converts raw images into continuous visual tokens, enabling efficient multimodal understanding.
-Crucially, the same Patch Embedding layer also encodes reference images for editing tasks, thereby superseding separate VAE-based tokenizers traditionally used [16, 43] and further enhancing inference efficiency.
-At its core, OneCAT integrates a Mixture-of-Experts (MoE) architecture.
-This MoE comprises three specialized feed-forward network (FFN) experts: one dedicated to processing text tokens for language comprehension, another designed for continuous visual tokens to facilitate multimodal understanding, and a third optimized for generating discrete visual tokens during image synthesis.
-All QKV and attention layers are shared across modalities and tasks, which promotes significant parameter efficiency and robust cross-modal alignment for instruction-following.
-For text generation, OneCAT adheres to the conventional Next-Token Prediction paradigm, leveraging the well-established capabilities of autoregressive language modeling.
-In parallel, for visual generation, it innovatively employs the Next-Scale Prediction paradigm [66, 24].
-This mechanism generates images in a coarse-to-fine, hierarchical manner, progressively predicting visual tokens from the lowest to the highest resolution scale, thereby achieving high-quality visual outputs.
-3.1 Architecture
-Our OneCAT model is initialized from the pre-trained Qwen2.5 LLM [82], leveraging its strong language modeling capabilities as a foundation.
-To construct our MoE architecture, we replicate the original FFN layer from each Qwen2.5 transformer block to form three distinct experts: a Text FFN (i.e., Text.
-FFN), a Visual Understanding FFN (i.e., Und.
-FFN), and a Visual Generation FFN (i.e., Gen.
-FFN).
-OneCAT employs a deterministic hard routing mechanism, where tokens are assigned to a specific expert based on their modality and the task at hand.
-The core functionality for each task is handled as follows:
+Figure 3: Inference pipeline of OneCAT, a decoder-only autoregressive unified model that seamlessly supports multimodal understanding, image generation and image editing.
+As depicted in Fig.
+3, OneCAT employs a pure decoder-only architecture and eliminates the need for additional vision encoder or tokenizer during inference.
+This streamlined design significantly simplifies the model structure and reduces computational cost.
+Unlike existing unified MLLMs [15, 21] using semantic encoders like ViTs for understanding, we follow recent encoder-free MLLMs [64] and use a lightweight Patch Embedding layer to convert raw images into continuous visual tokens for efficient and lossless processing.
+The same Patch Embedding layer is also used to process reference images for editing tasks, thereby superseding separate VAE tokenizer used [21, 55] and further enhancing inference efficiency.
+For text generation, OneCAT adheres to the Next-Token Prediction paradigm.
+For visual generation, it innovatively employs the Next-Scale Prediction [89], where images are generated in a hierarchical coarse-to-fine manner to progressively from the lowest to the highest resolution scales.
+Refer to Appendix A for more preliminaries for next-scale prediction.
+3.1 Modality-MoE
+OneCAT proposes a Modality-MoE architecture with three specialized feed-forward network (FFN) experts: a Text.
+FFN designed for text tokens for language comprehension, a Und.
+FFN designed for continuous visual tokens for visual understanding, and a Gen.
+FFN for discrete visual tokens for image generation.
+We employ a hard routing mechanism that assigns tokens to a specific expert based on their modality and the task at hand.
+Other QKV and attention layers are shared to promote parameter efficiency and robust cross-modal alignment for instruction-following.
+OneCAT is initialized from the pre-trained Qwen2.5 LLM [105] to exploit its strong ability in language modeling.
+To construct the proposed Modality-MoE, we replicate the FFN layer from each Qwen2.5 transformer block to form three distinct experts.
+The core functionality for each multimodal task is handled as below.
 Multimodal Understanding.
-To process images for understanding tasks, we employ a simple yet effective patch embedding layer that converts raw images into a sequence of continuous visual tokens.
-This layer consists of a $\times14$ convolution for image patchifying, a $\times2$ pixel unshuffle operation for visual token compression, and a two-layer Multilayer Perceptron (MLP) for projecting the visual features to match the LLM’s hidden state dimension.
-These continuous visual tokens are exclusively routed to the Und.
-FFN, while the text tokens for instruction are routed to the Text FFN.
+We employ a simple yet effective patch embedding layer to convert raw images into continuous visual tokens.
+This layer consists of a $\times14$ convolution for image patchifying, a $\times2$ pixel unshuffle layer, and a two-layer Multilayer Perceptron (MLP) for projecting the visual features to match the LLM’s hidden state dimension.
 Text-to-Image Generation.
-For image generation, we leverage a pre-trained multi-scale VAE model from Infinity [24] to map images between pixel space and latent space.
-This VAE operates with a downsampling ratio of 16 and a latent channel size of 32, and incorporates a bitwise quantizer [94] to enlarge the vocabulary.
-During training, the image tokenizer transforms the target images into a sequence of multi-scale discrete visual tokens to serve as ground-truth and input of LLM for teacher-forcing training, which are processed by the Gen.
+We leverage a pre-trained multi-scale VAE model from Infinity [31] to map images between the pixel and latent spaces.
+This VAE operates with a downsampling ratio of 16 and a latent channel size of 32, and uses a bitwise quantizer [118] to enlarge the vocabulary.
+During training, the VAE tokenizer encodes target images into multi-scale discrete visual tokens to serve as ground-truth.
+During inference, this tokenizer is not required, and only the detokenizer is used to reconstruct the image from generated multi-scale visual tokens.
+Image Editing.
+OneCAT seamlessly supports image editing task by leveraging a reference image and instruction input.
+The reference image is processed by the patch embedding layer, and the resulting continuous visual tokens are also routed to the Und.
+FFN to serve as the visual condition.
+The patch embedding layer provides a near-lossless representation of the reference image, and allows the LLM’s shallower layers to obtain low-level features for pixel-level consistency, while the deeper layers to extract high-level features for semantical comprehension.
+Guided by this hierarchical visual context, LLM predicts new discrete visual tokens autoregressively.
+This design enables powerful conditional generation without any architectural modifications, showing the versatility of our unified decoder-only design.
+3.2 Scale-Aware Adapter for Hierarchical Generation
+The tokens produced by the multi-scale VAE are inherently hierarchical (as shown in Appendix D).
+Specifically, lower-scale tokens mainly encode low-frequency global information such as color, illumination, and coarse structure, while higher-scale tokens capture high-frequency details including fine textures and intricate patterns.
+Processing these functionally divergent tokens equally with a shared Gen.
+FFN layer is not optimal and could limit the representation capacity.
+To address this, we introduce the Scale-Aware Adapter (SAA), a novel architectural component integrated with the Gen.
 FFN.
-Critically, during inference, the tokenizer is not required; only the detokenizer is needed to reconstruct the final image from the generated multi-scale visual tokens.
-The conditional text tokens are also routed to the Text FFN.
-Image Editing.
-OneCAT seamlessly supports image editing task by conditioning the visual generation process on a reference image and edit-instruction.
-The reference image is processed through the patch embedding layer, and the resulting continuous visual tokens are routed to the Und.
-FFN to serve as the visual condition, while the text tokens for instruction are routed to the Text FFN.
-The patch embedding layer provides a near-lossless representation of the reference image, which allows the LLM’s shallower layers to access low-level features for pixel-level consistency, while the deeper layers extract high-level features for semantical comprehension.
-Guided by this rich, hierarchical visual context, the model then autoregressively predicts new discrete visual tokens, which are handled by the Gen.
+The SAA comprises a set of parallel modules that serve as skip connections over each linear layer of the Gen.
 FFN.
-This design enables powerful conditional generation without requiring any architectural modifications, showcasing the versatility of our unified decoder-only design.
-3.2 Scale-Aware Adapter for Hierarchical Generation
-While the Gen.
-FFN is capable of processing discrete visual tokens, the tokens produced by the multi-scale VAE are inherently hierarchical.
-A standard FFN would treat tokens for different scales equally, ignoring this crucial structural information.
-To address this and enable more granular control over the visual generation process, we introduce the Scale-Aware Adapter (SAA), a novel architectural component integrated with the Gen.
-FFN.
-Our design is motivated by the principle that different scales in the multi-scale VAE govern distinct aspects for image generation.
-Specifically, tokens from lower scales predominantly encode low-frequency global information, such as color, illumination, and coarse structure.
-Those from higher scales, conversely, capture high-frequency details including fine textures and intricate patterns.
-Processing these functionally divergent tokens with a shared Gen.
-FFN layer limits representational capacity and is thus suboptimal.
-The Scale-Aware Adapter (SAA) comprises a set of parallel modules that serve as skip connections over the Gen.
-FFN.
-Each module is dedicated to processing tokens from a specific scale of the multi-scale VAE, with the total number of adapters matching the number of VAE scales.
-During inference, discrete visual tokens are routed to their corresponding scale-specific adapter based on the scale index.
-To ensure parameter efficiency, each adapter is constructed using a low-rank decomposition (rank rr= 64), inspired by the LoRA [26].
-However, unlike LoRA which is typically used for fine-tuning, the SAA modules are trained jointly end-to-end as permanent components of the LLM.
+Each SAA is dedicated to processing tokens from a specific scale, with the total number of SAA of a Gen.
+FFN matching the number of VAE scales.
+During inference, discrete visual tokens are routed to corresponding scale-specific adapters based on scale indices.
+To ensure parameter efficiency, each adapter is achieved using a low-rank decomposition (rank rr= 64), inspired by LoRA [33].
+However, unlike LoRA that is typically used for fine-tuning, the SAA modules are jointly trained in an end-to-end manner as permanent components of the LLM.
 3.3 Multimodal versatile attention mechanism
-We leverage a multimodal versatile attention mechanism based on PyTorch FlexAttention [57] to empower OneCAT with the ability to process diverse modalities and tasks in a flexible and adaptive manner.
-As illustrated in Fig.
-4, text tokens $T$ are processed using causal attention, ensuring autoregressive generation; Continuous visual tokens $U$ are processed via full attention, allowing each token to attend to all others in the sequence.
-Multiscale discrete visual tokens $G_{i}$ (where $i$ denotes the scale index) are processed via block causal attention, tokens within the same scale can attend to each other freely, while attention across scales follows a causal attention.
+We employ a multimodal versatile attention mechanism based on FlexAttention [76] to enable flexible processing of diverse modalities and tasks within a single LLM.
+As shown in Fig.
+4, text tokens $T$ use causal attention for autoregressive generation, while continuous visual tokens $U$ apply full attention for global interaction.
+Multiscale discrete visual tokens $G$ use block causal attention.
+Tokens within the same scale attend to each other freely, while attention across scales follows a causal attention.
 Figure 4: Multimodal versatile attention mechanism.
@@
 4 Model Training Pipeline
-The training pipeline is divided into three stages and the instruction of the training recipe is shown in Tab. 1.
+As shown in Fig. 5 and Table 1, we employ a three-stage training pipeline for OneCAT.
 Figure 5: Overview of the training pipeline.
@@
 4.1.1 Stage 1-1: Teacher Training
-Rather than employing an off-the-shelf MLLM as teacher (e.g., Qwen2.5-VL [3]), we construct a custom teacher model to ensure parameter consistency between the LLM backbones of the teacher and student models, thereby improving distillation efficiency.
-Specifically, the teacher is built by connecting a pre-trained vision encoder (InternViT [14]) and a LLM (Qwen2.5 [82]) with a two-layer MLP as connector.
-During this sub-stage, we freeze the ViT and LLM, and exclusively train the MLP connector on a curated small-scale dataset of image-to-text caption pairs using the standard NTP loss.
+Instead of using an off-the-shelf MLLM as teacher (e.g., Qwen2.5-VL [5]), we customize a teacher to ensure parameter consistency between the LLM backbones of the teacher and student, and thus, improve distillation stability and efficiency (refer to Fig.
+8).
+Specifically, the teacher is built by connecting a pre-trained ViT (InternViT [19]) and a LLM (Qwen2.5 [105]) with a two-layer MLP.
+We freeze both ViT and LLM and only train the MLP connector on a small-scale image-text caption dataset with NTP loss.
+This alignment training endows the custom MLLM teacher with strong visual perception ability.
 4.1.2 Stage 1-2: Expert Pretraining
-With the teacher model prepared, we proceed to train the OneCAT model.
-We keep the QKV Projection, Attention, and Text FFN frozen, and selectively optimize the task-specific modules: the Und.
-FFN and Patch Embedding Layer for multimodal understanding, and the Gen.
-FFN for text-to-image generation.
+We then train OneCAT based on the teacher model.
+We freeze the QKV, Attention, and Text FFN, and optimize the task-specific modules: the Und.
+FFN and Patch Embedding layer for visual understanding, and the Gen.
+FFN for visual generation.
 Hyperparameter / Config Stage 1-1 Stage 1-2 Stage 2 Stage 3
@@
 Understanding Distillation: We optimize the Und.
-FFN on a large-scale dataset of image-to-text caption pairs to acquire fundamental visual knowledge.
-Based on the specifically designed teacher, the training objective is a combination of the NTP loss ($\mathcal{L}_{\text{NTP}}$) and a distillation loss ($\mathcal{L}_{\text{Distill}}$).
-Specifically, $\mathcal{L}_{\text{NTP}}$ is the standard cross-entropy loss for autoregressive text generation.
-For distillation, instead of matching the final output logits, we align the student’s internal hidden states with those of the teacher model through deep feature-level matching.
-This strategy enables the student to not only mimic the teacher’s final prediction but also its intermediate computational patterns across all token positions (including both visual and text tokens) for better visual knowledge transfer.
-The distillation loss $\mathcal{L}_{\text{Distill}}$ is formulated as the sum of MSE losses between the hidden states of the student and teacher models over all $N$ transformer layers:
-$\mathcal{L}_{\text{Distill}}=\sum_{n=1}^{N}\text{MSE}(\mathbf{h}_{S}^{(n)},\mathbf{h}_{T}^{(n)}),$ (1)
-where $\mathbf{h}_{S}^{(n)}$ and $\mathbf{h}_{T}^{(n)}$ represent the hidden state outputs from the $n$-th transformer block of the student and teacher models, respectively.
-The final objective is thus:
-$\mathcal{L}_{\text{Und}}=\mathcal{L}_{\text{NTP}}+\lambda\mathcal{L}_{\text{Distill}},$ (2)
-where $\lambda=0.2$ is a balancing hyperparameter.
+FFN on a large-scale dataset of image-to-text pairs.
+The training objective is a combination of the NTP loss ($\mathcal{L}_{\text{NTP}}$) and a distillation loss ($\mathcal{L}_{\text{Distill}}$).
+Specifically, $\mathcal{L}_{\text{NTP}}$ is the cross-entropy loss for text token generation.
+For distillation, instead of matching the final output logits, we align the student’s internal hidden states with those of the teacher model through deep feature-level matching across each transformer layer.
+This enables the student to not only mimic the teacher’s final text prediction but also its internal computational patterns across all token (including both visual and text tokens) for better visual knowledge transfer.
+The distillation loss is formulated as $\mathcal{L}_{\text{Distill}}=\sum_{n=1}^{N}\text{MSE}(\mathbf{h}_{S}^{(n)},\mathbf{h}_{T}^{(n)})$, where $\mathbf{h}_{S}^{(n)}$ and $\mathbf{h}_{T}^{(n)}$ represent the hidden state outputs from the $n$-th transformer block of the student and teacher models, respectively.
+The final objective is $\mathcal{L}_{\text{Und}}=\mathcal{L}_{\text{NTP}}+\lambda\mathcal{L}_{\text{Distill}}$, where $\lambda=0.02$ is a balancing hyperparameter.
 Throughout distillation, all models process images at a fixed resolution of $448\times 448$ to balance computational load and the granularity of visual features.
 Generation Pretraining: In parallel, we optimize the Gen.
-FFN on a delicate text-to-image generation dataset to enable the model to learn the spatial relationships and dependencies between multi-scale discrete visual tokens.
-We adopt the cross-entropy loss for next-scale prediction [66, 24] and the output image resolution is fixed to $\times256$ in this stage.
+FFN on a delicate text-to-image generation dataset to enable the LLM to learn the spatial relationships and cross-scale dependencies.
+We also adopt the cross-entropy loss for next-scale prediction [89] and the output image resolution is fixed to $\times256$.
 4.2 Stage-2: Unified Mid-Training
-In the second stage, we unfreeze the entire model to achieve unified mid-training across multiple tasks (i.e., image-to-text, text-to-image, image editing, and text-only dialogues).
-At this stage, we incorporate the proposed scale-aware adapter (SAA), which is optimized with other modules of LLM together to extract scale-specific representation for enhanced image generation.
-We introduce native resolution strategy for both understanding and generation in this stage.
-For multimodal understanding, the model is trained to process images at their original resolutions, thereby preserving fine-grained details and eliminating information loss.
+In Stage-2, we unfreeze the entire model to perform unified mid-training across multiple tasks (i.e., image-to-text, text-to-image, image editing, and text-only dialogues).
+We incorporate the proposed scale-aware adapter in this stage, which is optimized with other modules together to extract scale-specific representation for enhanced image generation quality.
+We also introduce native resolution strategy in this stage.
+For visual understanding, the model is trained to process images at their original resolutions, thereby preserving fine-grained details and reducing information loss.
 Additionally, a thumbnail of resolution $\times448$ is included to provide global visual context.
-For visual generation, the model is trained to generate images with different resolution and aspect ratios where the side lengths dynamically sampled from a range of 288 to 864 pixels, which significantly enhancing its generative versatility and real-world applicability.
+For visual generation, we train the model with dynamical resolution and aspect ratios where the side lengths sampled from a range of 288 to 864 pixels, which enhancing its generation versatility and real-world applicability.
 4.3 Stage-3: Unified Supervised Fine-tuning
 The final stage involves unified supervised fine-tuning (SFT) using a curated dateset of higher-quality data to enhance instruction-following and visual generation quality.
-The native resolution strategy was continued, with the size of generated image expanded to support side lengths between 288 to 1728 pixels, enabling higher-resolution results.
+The native resolution strategy was continued, with the size of generated image expanded to support side lengths between 288 to 1776 pixels, enabling higher-resolution results.
 5 Data Setup
@@
 This dataset is collected from two primary sources: (1) Public Available Image-Text Caption Pairs: We incorporate several publicly available, high-quality image-caption datasets, including Recap-DataComp-1B [52], Capsfusion [110], Detailed-Caption [53], SA1B-Dense-Caption [20], and Moondream2-COYO-5M-Captions [40].
-(2) Re-captioned Image Datasets: We generate new image-text pairs by re-captioning large-scale public image collections.
+(2) Re-captioned Image Datasets: We generate new image-text pairs by re-captioning large-scale public image collections using Qwen2-VL [92].
 The source image datasets for this process include COYO700M [8], CC12M [9], CC3M [80], LAION-400M [79], and Zeor250M [101].
-From this dataset, we randomly select a subset of 10 million samples to train the custom teacher.
+From this large-scale dataset, we randomly sample a small-scale subset of 10 million samples to train the custom teacher.
 For image generation, we construct a dataset of 52 million text-to-image samples after a rigorous filtering process to remove samples with low resolution or poor aesthetic scores.
-This collection consists of 1 million class-labeled images from ImageNet-1k [17], 20 million pairs from re-captioned public collections (i.e., COYO700M [6], LAION-400M [59] and CC12M [7]), and 30 million in-house synthetic images generated by FLUX.
+This collection consists of 1 million class-labeled images from ImageNet-1k [22], 20 million pairs from public collections (i.e., COYO700M [8], LAION-400M [79] and CC12M [9]), and 30 million in-house synthetic images generated by FLUX.
 The overall training token ratio across multimodal understanding and visual generation samples in Stage-I is approximately 8:1.
-Stage-2: In the unified mid-training, for multimodal understanding we leverage an in-house dataset of 70 million visual instruction samples.
-This dataset is specifically curated to be highly diverse tasks, including general VQA, detailed image captioning, OCR, multimodal reasoning(i.e., STEM problem-solving), knowledge, and visual grounding.
+Stage-2: In the unified mid-training, for multimodal understanding we leverage an curated dataset of 70 million visual instruction samples.
+This dataset is specifically curated to be highly diverse tasks, including general VQA, detailed image captioning, OCR, multimodal reasoning(i.e., STEM problem-solving), knowledge, and visual grounding, which are sourced from Detailed-Caption [53], ALLaVA [10], ShareGPT4V [14], SA1B-Dense-Caption [20], WIT [85], pdfa-eng-wds [1], UReader [106], DVQA [69], OCR-VQA [71], WebSRC [16], GQA [37], visual-genome [44], GRIT [74], and other in-house synthetic visual instruction data.
 For visual generation, we supplement the text-to-image samples of Stage-1 with a additional collection of 8 million image editing samples, resulting a total of 60 million visual generation samples.
@@
 To ensure a strong focus on visual generation in Stage-II, we oversample the visual generation data, resulting a final training token ratio of approximately 1 :2 :6 across text-only, multimodal understanding, and visual generation tasks, respectively.
-OneCAT-1.5B OneCAT-3B
-Base Model Qwen2.5-1.5B-instruct Qwen2.5-3B-instruct
-Active Parameters 1.5B 3B
-Total Parameters 4.5B 9B
-Understanding Distillation
-Teacher ViT InternViT-300M
-Teacher LLM Qwen2.5-1.5B-instruct Qwen2.5-3B-instruct
-Table 2: Model configurations for the two variants of OneCAT.
-Stage-3: In the SFT stage, for multimodal understanding and text-only instruction, we construct a high-quality dataset of 13 million samples, combining 10 million from public MAmmoTH-VL dataset [22] with 3 million of our in-house synthetic instruction samples to further improve the reasoning abilities.
-For visual generation, we utilize a total of 3 million samples, aggregated from UniWorld [46], blip3o-60k [8], ShareGPT-4o-Image [10], and additional in-house synthetic data generated by GPT-4o and FLUX.
-The overall training token ratio across text-only, multimodal understanding, and visual generation is approximately 1 :5 :6.
+Table 2: Model configurations for the two variants of OneCAT. A params indicates the activated parameters and T params indicates the total parameters.
+Model Recipe Understanding Distillation
+Base Model A Params T Params Teacher ViT Teacher LLM
+OneCAT-1.5B Qwen2.5-1.5B-instruct 1.5B 4.5B InternViT-300M Qwen2.5-1.5B-instruct
+OneCAT-3B Qwen2.5-3B-instruct 3B 9B InternViT-300M Qwen2.5-3B-instruct
+Stage-3: In the SFT stage, for multimodal understanding and text-only instruction, we construct a high-quality dataset of 13 million samples.
+This dataset comprises 10 million filtered samples from MAmmoTH-VL dataset [29] and 3 million samples from other open-source datasets AI2D [43], OKVQA [66], VQAv2 [28], ART500K [65], ScienceQA [78], GQA [37], CLEVR-Math [59], COCO-ReM [84], TallyQA [2], Docmatix [46], DVQA [69], DreamSim [26], ShareGPT4o [18].
+For visual generation, we utilize a total of 3 million samples, aggregated from UniWorld [58], BLIP3o-60k [11], ShareGPT-4o-Image [13], and additional synthetic data generated by GPT-4o [39] and FLUX [45] using the partial prompts from JourneyDB [86].
+The overall training token ratio across text-only, multimodal understanding, and visual generation for unified sft is approximately 1 :5 :6.
 6 Implementation details
-Model Configurations: We conduct our experiments on two model variants, OneCAT-1.5B and OneCAT-3B.
-The OneCAT-1.5B model is based on Qwen2.5-1.5B-instruct [3] and contains 1.5B active parameters (4.5B total).
-The OneCAT-3B model is built upon Qwen2.5-3B-instruct [3] and utilizes 3B active parameters (9B total).
-For understanding distillation, the ViT of teacher model is InternViT-300M [14] for both two variants and the LLM of teacher model is Qwen2.5-1.5B-instruct and Qwen2.5-3B-instruct for OneCAT-1.5B and OneCAT-3B, respectively, to align with the LLM backbone of our OneCAT.
-Detailed instruction of the model configurations is shown in Tab.
-2.
+Model Configurations: We conduct experiments on two model variants, OneCAT-1.5B and OneCAT-3B, which is initialized from Qwen2.5-1.5B-instruct and Qwen2.5-3B-instruct, respectively.
+The total parameters count of OneCAT-3B is 9B, but the activated parameters count for each token is 3B during forward process.
+Refer to Table 2 for more details.
 Table 3: Performance comparison across multiple multimodal understanding benchmarks.
@@
 Encoder-free Understanding Only Models
-Mono-InternVL [51] 1.8B / 72.6 73.7 - - 59.5 68.6
+Mono-InternVL [64] 1.8B / 72.6 73.7 43.0 83.0 59.5 68.6
 EvE [23] 7B / 56.8 59.1 - - 62.6 61.0
@@
 Similar to [56], we then employ Global Batch Reduced Loss by dividing each micro-batch loss by the averaged token count per process, $\tfrac{T_{global}}{N}$, which can be shown that the final synchronized gradient for the subsequent optimizer step is mathematically equivalent to the gradient of the global batch loss, enabling training with unbiased gradients.
+Classifier-free guidance (CFG).
+We follow previous works [13, 21] to use CFG for enhanced visual generation quality.
+For training, we randomly drop conditional text and reference image tokens.
+For inference, we combines conditional and unconditional predicted logits.
+For more details, please refer to Appendix B.
 Table 4: Performance comparison across multiple multimodal understanding benchmarks.
@@
 7.1 Multimodal understanding.
-We evaluate our model on 12 public Multimodal understanding benchmarks spanning diverse capabilities: MMbench [49], MME [84], MMMU [88], MM-Vet [87], and SEED [36] assess general multimodal perception and reasoning.
+We evaluate OneCAT on several multimodal understanding benchmarks: MMbench [62], MME [108], MMMU [112], MM-Vet [111], and SEED [48] assess general multimodal perception and reasoning.
 MathVista [63] focuses on mathematical reasoning.
@@
 This highlights the effectiveness of our model’s ability to condition its generation on fine-grained visual cues from a reference image.
-Figures 6 and 7 present qualitative comparisons for the text-to-image and image-editing tasks, respectively.
-OneCAT-3B exhibits leading instruction-following and world-understanding capabilities among open-source models.
-For example, in the fourth row of Figure 6, it is the only open-source model that correctly generates exactly four characters in a pixel-art style; in contrast, BAGEL-7B produces five characters, and Janus-Pro-7B fails to render pixel art.
-Similarly, in the last row of Figure 7, only OneCAT-3B produces an image with the correct lighting conditions.
-In summary, the strong instruction-following capability in both text-to-image generation and image editing tasks highlights the effectiveness and versatility of our architectural design.
+7.3 Qualitative results.
+We present qualitative comparisons for the text-to-image and image-editing tasks in Fig.
+6 and 7, respectively.
+OneCAT-3B exhibits leading instruction-following and world-understanding capabilities.
+We also present more visual generation and understanding showcases in Appendix E.
 Table 6: Performance comparison on the DPG-Bench [34] benchmark. Best in bold, second best is underlined.
@@
 In addition to its strong performance, our model’s architectural design also yields significant improvements in inference efficiency for handling both high-resolution image input and output.
-We evaluated the efficiency of OneCAT-3B in two distinct phases: prefilling and generation.
-In the prefilling phase, we measure the Time to the First Token (TTFT) to assess the computational cost of processing the input.
+The left section of Tab.
+8 compares the inference efficiency of OneCAT and Qwen2.5-VL for multimodal understanding.
+Benefiting from our pure decoder-only architecture, which removes the external ViT encoder, OneCAT achieves significantly faster prefilling, reducing first-token latency by up to 61% compared to Qwen2.5-VL on high-resolution inputs.
+As shown in the right section of Tab.
+8, OneCAT exhibits a substantial speed advantage over the diffusion-based BAGEL in image generation.
+When producing a $1024\times 1024$ image, OneCAT requires only 2.85s for T2I generation and 4.61s for image editing—approximately $10\times$ faster than BAGEL.
+This improvement stems from two key elements: our multi-scale autoregressive mechanism within the LLM, and a VAE tokenizer-free design for editing that reduces both encoding and decoding overhead.
+Table 8: Efficiency comparison for understanding (Left) and generation (Right), tested on one NVIDIA H800.
+Left: We report the time to first token (TTFT).
+The number of input text tokens are fixed to 24.
+256* is the number of visual tokens of thumbnail.
+Right: We report total inference time for Text-to-Image (T2I) and Image-Editing.
+Model Resolution of Input Image #Tokens Visual TTFT(s)
+Qwen2.5-VL-3B $768\times 768$ 731 0.135
+[…]
+Qwen2.5-VL-3B $1792\times 1792$ 4098 0.583
+[…]
+Model Resolution of Output Image T2I Infer. Time (s) Edit Infer. Time (s)
+BAGEL-7B $512\times 512$ 8.76 13.45
+[…]
+BAGEL-7B $1024\times 1024$ 26.29 46.44
+[…]
+7.5 Ablation Study
+7.5.1 Ablation Study on Understanding Distillation (Und. Distil.) Stage
+We ablate each module in the Und.
+Distil.
+stage using the OneCAT-1.5B.
+Each experiment uses 8B sampled tokens ($\sim$10M image-text pairs from Stage-1 dataset) to optimize the Und.
+FFN, followed by a simplified SFT using LLaVA-665k [60] dataset.
+On Different Distillation Strategies.
 As shown in Tab.
-8, thanks to our pure decoder-only architecture that eliminates the need for a separate ViT encoder, OneCAT-3B demonstrates a substantial efficiency advantage, particularly when handling high-resolution images.
-Compared to QwenVL-3B, as the image resolution increases from $768\times 768$ to $1792\times 1792$, the reduction in TTFT for OneCAT-3B grows sharply from 50.4% to 61.4%.
-This provides strong evidence for the effectiveness of our design, especially for high-resolution inputs.
-In the generation phase, we evaluated the total inference time for text-to-image (T2I) generation and image editing.
-As detailed in Tab.
-9, OneCAT-3B’s inference speed far surpasses that of the strong diffusion-based baseline, BAGEL-7B.
-For instance, when generating a $1024\times 1024$ resolution image, OneCAT-3B’s T2I and editing inference times are merely 2.85s and 4.61s, respectively, making it approximately $10\times$ faster than BAGEL-7B.
-This significant advantage stems from our pioneering multi-scale autoregressive generation mechanism inside LLM that enables parallel token generation, which drastically reduces the number of required decoding steps compared to iterative diffusion models.
-In summary, OneCAT achieves exceptional inference efficiency in both the prefilling and generation stages.
-This highlights the immense potential and practical value of our unified autoregressive framework for building efficient and powerful large multimodal models.
-Table 8: Efficiency comparison of OneCAT-3B and Qwen2.5-VL-3B.
-Models are tested based on one NVIDIA H800 GPU.
-We report the time to the first token (TTFT) to measure the reduction of computational cost of prefilling phase.
-256* denotes the number of visual tokens for thumbnail.
-Model Resolution of Input Image #Input Text Tokens #Input Visual Tokens TTFT(s) Reduction
-Qwen2.5-VL-3B $768\times 768$ 24 731 0.135 50.4%
-OneCAT-3B $768\times 768$ 24 731 $256^{*}$ 0.067
-Qwen2.5-VL-3B $1024\times 1024$ 24 1395 0.216 57.4%
-OneCAT-3B $1024\times 1024$ 24 1395 $256^{*}$ 0.092
-Qwen2.5-VL-3B $1792\times 1792$ 24 4098 0.583 61.4%
-OneCAT-3B $1792\times 1792$ 24 4098 $256^{*}$ 0.225
-Table 9: Generation efficiency comparison of OneCAT-3B and BAGEL.
-Models are tested based on one NVIDIA H800 GPU.
-We report total inference time to measure the Text-to-Image(T2I) and Image-Editing efficiency.
-Model Resolution of Generated Image T2I Infer. Time (s) Edit Infer. Time (s)
-BAGEL-7B $512\times 512$ 8.76 13.45
-OneCAT-3B $512\times 512$ 1.40 2.03
-BAGEL-7B $1024\times 1024$ 26.29 46.44
-OneCAT-3B $1024\times 1024$ 2.85 4.61
+9 and Fig.
+8(a), our proposed distillation approach—distilling all layers’ hidden states—significantly improves visual learning efficiency and multimodal performance.
+In addition, distilling only the last layer’s hidden states or logits also brings gains over the baseline, yet remains inferior to our full distillation approach.
+This highlights the necessity of full-layer distillation for effective visual knowledge transfer from the custom teacher MLLM.
+Notably, our strategy yields better performance than applying the distillation approaches of EvE [23] and VoRA [91] to our OneCAT-1.5B under the same setting, further validating the efficacy of our proposed method.
+Methods MMB MME-S MMVet SEED AI2D ChartQA TextVQA Avg.
+w/o distillation 42.0 1209 13.2 47.0 53.8 10.5 10.0 31.4
+distill last layer’s logits 44.8 1312 15.4 53.0 55.2 11.5 10.6 33.9
+distill last layer’s hidden states 46.0 1255 15.3 54.3 55.3 11.0 10.9 33.9
+[…]
+EvE [23] 44.5 1276 15.0 52.8 54.0 10.6 11.2 33.4
+VoRA [91] 41.5 1234 15.1 50.6 54.2 10.4 10.5 32.3
+Table 9: Effect of different distillation strategies.
+Methods Avg.
+w/o distillation 31.4
+[…]
+distill with Qwen2.5-VL teacher 33.7
+Table 10: Effect of different teachers. Avg. denotes the average score in Tab. 9.
+On Different Teachers.
+We then analyze the effect of different teacher models.
+As shown in Tab.
+10 and Fig.
+8(b), using the pretrained Qwen2.5-VL-1.5B as the teacher results in unstable training and lower distillation efficiency due to parameter misalignment between the teacher and student LLMs—especially for the frozen attention and QKV layers—highlighting the importance of our custom teacher design.
+Figure 8: Comparision of different distillation strategies and teachers for stage-1 training.
+7.5.2 Ablation Study on Unified Mid-Training (Stage-2)
+We initialize the model with OneCAT-1.5B from Stage-1 and perform a simplified unified mid-training to study (i) the effect of token ratio across tasks and (ii) the effect of SAA.
+All models then undergo a simplified SFT using LLaVA-665k [60] for understanding and BLIP3o-60k [11] for generation with a 1:1 token ratio.
+On Token Ratio.
+We sample 5B and 10B tokens for text-only (T) and multimodal understanding (U), and vary the number of trained tokens in visual generation (G) to study the impact of token ratio.
+As shown in Table 11, increasing the training tokens for G does not significantly affect U but improves G.
+To balance training cost and performance, we adopt a token ratio of 1:2:6 for Stage-2.
+Token Ratio Understanding Generation
+(T:U:G) MMB MME-S MMVet SEED AI2D ChartQA TextVQA Avg. GenEval DPG
+5B : 10B : 0B 61.6 1556 28.9 67.2 60.1 56.1 56.2 55.1 - -
+5B : 10B : 10B 62.7 1547 29.7 66.7 60.3 59.8 55.4 55.6 80.7 74.6
+[…]
+5B : 10B : 45B 62.1 1544 30.1 66.8 58.4 58.6 55.0 55.2 81.4 75.6
+Table 11: Effect of trained token ratio across text-only (T), understanding (U), and generation (G).
+GenEval DPG
+[…]
+w/o SAA 78.1 74.0
+Table 12: Effect of SAA.
+On SAA.
+We then remove the SAA during the stage-2 training for performance comparison, as shown in Table 12.
+Each model train 5B, 10B, 30B tokens for T, U, and G for Stage-2, respectively.
+We can see that removing SAA leads to an obvious performance drop.
+We also present the visualization of frequency properties of tokens from different scales in Appendix D to better understand the motivation of our SAA.
 8 Conclusion
@@
 Our results demonstrate the viability and advantages of a first-principles approach to multimodal modeling, offering a powerful new baseline for future research and applications in general-purpose multimodal intelligence.
+Appendix A Preliminary of Next-Scale Prediction
+A.1 Multiscale Tokenization
+Leveraging the inherent coarse-to-fine structure of natural images, VAR [89] introduces a multi-scale tokenizer that encodes an image into $K$ token scales $(R_{1},R_{2},\dots,R_{K})$.
+The resolution of each scale $R_{k}$, denoted as $(h_{k},w_{k})$, increases monotonically with the scale index $k$.
+Specifically, given a feature map $F$ extracted from an image with an image encoder, VAR defines these token scales recursively:
+$R_{1}$ $=\mathcal{Q}(\operatorname{interpolate}_{1}(F)),$ (1)
+$R_{2}$ $=\mathcal{Q}(\operatorname{interpolate}_{2}(F-\operatorname{interpolate}_{K}(R_{1}))),$ (2)
+$\vdots$
+$R_{k}$ $=\mathcal{Q}\Big(\operatorname{interpolate}_{k}(F-\sum_{i=1}^{k-1}\operatorname{interpolate}_{K}(R_{i}))\Big),$ (3)
+$\vdots$
+$R_{K}$ $=\mathcal{Q}\Big(F-\sum_{i=1}^{K-1}\operatorname{interpolate}_{K}(R_{i})\Big),$ (4)
+where $\operatorname{interpolate}_{i}$ is an operator that resizes its input to the resolution $(h_{i},w_{i})$, and $\mathcal{Q}$ is the quantization operator.
+For a given 3D feature map $x\in\mathbb{R}^{d\times h\times w}$, we implement $\mathcal{Q}$ using Binary Spherical Quantization (BSQ) [118], following Han et al.
+[31].
+The quantization is applied to each spatial feature vector $x_{ij}\in\mathbb{R}^{d}$ as:
+$\mathcal{Q}(x_{ij})=\frac{1}{\sqrt{d}}\operatorname{sign}\left(\frac{x_{ij}}{\|x_{ij}\|_{2}}\right).$ (5)
+A.2 Visual Auto-Regressive Training
+The premise for generation is that the feature map $F$ can be well approximated by summing all scales upsampled to the final resolution: $F\approx\sum_{i=1}^{K}\operatorname{interpolate}_{K}(R_{i})$.
+It therefore suffices to generate the sequence of scales $R_{1:K}$ to synthesize an image.
+To achieve this, VAR models the joint distribution over the scales auto-regressively, factorizing the log-likelihood as:
+$\log p_{\theta}(R_{1:K})=\sum_{k=1}^{K}\log p_{\theta}(R_{k}\mid R_{1:k-1}).$ (6)
+The model, with parameters $\theta$, is trained to maximize this log-likelihood by learning to predict the current scale $R_{k}$ conditioned on all preceding scales $R_{1:k-1}$.
+To enable efficient parallel decoding, VAR assumes that all tokens within the current scale $R_{k}$ are conditionally independent given $R_{1:k-1}$.
+However, this conditional independence assumption, coupled with imperfect model training, can lead to error propagation: mistakes in generating early-stage scales ($R_{1},\dots,R_{k-1}$) are amplified when generating subsequent, higher-resolution scales $R_{k}$.
+To mitigate this issue, Han et al.
+[31] proposed Bitwise Self-Correction.
+This technique involves training the model on corrupted versions of the conditioning scales $R_{1:k-1}$, thereby teaching it to generate the correct $R_{k}$ even when the preceding scales are imperfect.
+This robustifies the model against its own generation errors during inference.
+A.3 Predefined Scale Schedules
+We follows Han et al.
+[31] and Tian et al.
+[89] to establish a set of predefined scale schedules, thus ensuring efficient training across images with varying aspect ratios.
+As detailed in Table 13, for each target aspect ratio $r$, we define a specific schedule as a sequence of $K$ resolution tuples: $\{(h_{1}^{r},w_{1}^{r}),\dots,(h_{K}^{r},w_{K}^{r})\}$.
+These schedules are designed based on two fundamental principles: 1.Aspect Ratio Consistency: Each tuple $(h_{k}^{r},w_{k}^{r})$ within a schedule maintains an aspect ratio that is approximately equal to the target ratio $r$, especially at larger scales.
+2.Consistent Area Across Scales: For any given scale level $k$, the image area, calculated as $h_{k}^{r}\times w_{k}^{r}$, is kept roughly constant across different aspect ratio schedules.
+This standardization ensures that the training sequence lengths are similar for various aspect ratios, thereby improving overall training efficiency.
+During the inference stage, these predefined schedules enable the model to generate high-quality images covering a wide range of aspect ratios.
+Table 13: Predefined scale schedules $\{(h_{1}^{r},w_{1}^{r}),\dots,(h_{K}^{r},w_{K}^{r})\}$ for different aspect ratios.
+Following Han et al.
+[31], OneCAT utilizes $K=13$ scales to generate the highest resolution image, such as a $1024\times 1024$ image for the 1:1 aspect ratio, while lower-resolution images like $512\times 512$ can be produced by truncating the schedule to $K=10$.
+Aspect Ratio Resolution Scale Schedule
+1.000 (1:1) $1024\times 1024$ (1,1) (2,2) (4,4) (6,6) (8,8) (12,12) (16,16) (20,20) (24,24) (32,32) (40,40) (48,48) (64,64)
+0.800 (4:5) $896\times 1120$ (1,1) (2,2) (3,3) (4,5) (8,10) (12,15) (16,20) (20,25) (24,30) (28,35) (36,45) (44,55) (56,70)
+1.250 (5:4) $1120\times 896$ (1,1) (2,2) (3,3) (5,4) (10,8) (15,12) (20,16) (25,20) (30,24) (35,28) (45,36) (55,44) (70,56)
+0.750 (3:4) $864\times 1152$ (1,1) (2,2) (3,4) (6,8) (9,12) (12,16) (15,20) (18,24) (21,28) (27,36) (36,48) (45,60) (54,72)
+1.333 (4:3) $1152\times 864$ (1,1) (2,2) (4,3) (8,6) (12,9) (16,12) (20,15) (24,18) (28,21) (36,27) (48,36) (60,45) (72,54)
+0.666 (2:3) $832\times 1248$ (1,1) (2,2) (2,3) (4,6) (6,9) (10,15) (14,21) (18,27) (22,33) (26,39) (32,48) (42,63) (52,78)
+1.500 (3:2) $1248\times 832$ (1,1) (2,2) (3,2) (6,4) (9,6) (15,10) (21,14) (27,18) (33,22) (39,26) (48,32) (63,42) (78,52)
+0.571 (4:7) $768\times 1344$ (1,1) (2,2) (3,3) (4,7) (6,11) (8,14) (12,21) (16,28) (20,35) (24,42) (32,56) (40,70) (48,84)
+1.750 (7:4) $1344\times 768$ (1,1) (2,2) (3,3) (7,4) (11,6) (14,8) (21,12) (28,16) (35,20) (42,24) (56,32) (70,40) (84,48)
+0.500 (1:2) $720\times 1440$ (1,1) (2,2) (2,4) (3,6) (5,10) (8,16) (11,22) (15,30) (19,38) (23,46) (30,60) (37,74) (45,90)
+2.000 (2:1) $1440\times 720$ (1,1) (2,2) (4,2) (6,3) (10,5) (16,8) (22,11) (30,15) (38,19) (46,23) (60,30) (74,37) (90,45)
+0.400 (2:5) $640\times 1600$ (1,1) (2,2) (2,5) (4,10) (6,15) (8,20) (10,25) (12,30) (16,40) (20,50) (26,65) (32,80) (40,100)
+2.500 (5:2) $1600\times 640$ (1,1) (2,2) (5,2) (10,4) (15,6) (20,8) (25,10) (30,12) (40,16) (50,20) (65,26) (80,32) (100,40)
+0.333 (1:3) $592\times 1776$ (1,1) (2,2) (2,6) (3,9) (5,15) (7,21) (9,27) (12,36) (15,45) (18,54) (24,72) (30,90) (37,111)
+3.000 (3:1) $1776\times 592$ (1,1) (2,2) (6,2) (9,3) (15,5) (21,7) (27,9) (36,12) (45,15) (54,18) (72,24) (90,30) (111,37)
+Appendix B Details of Class-Free Guidance
+We follow previous works [13, 21] to use CFG for enhanced visual generation quality.
+For training, we randomly drop tokens of conditional text and reference image with probabilities 0.1.
+For inference, we combine conditional and unconditional predicted logits to produce outputs.
+Specifically, for text-to-image generation, the final logits $\textbf{L}_{\text{final}}$ are computed as a linear combination of the conditional logits $\textbf{L}_{t}$ (with text input) and unconditional logits $\textbf{L}_{\emptyset}$ (without text input):
+$\textbf{L}_{\text{final}}=\lambda_{t}\cdot\textbf{L}_{t}+(1-\lambda_{t})\cdot\textbf{L}_{\emptyset}$ (7)
+where $\lambda_{t}$ is the text guidance scale controlling the influence of the text condition.
+For image editing tasks, which involve both textual and reference image conditions, we employ a dual-guidance mechanism.
+Let $\textbf{L}_{t,i}$ denote the logits with both text and reference image conditions, $\textbf{L}_{t}$ the logits with text condition only, and $\textbf{L}_{\emptyset}$ the logits without any conditions.
+The refined logits $\textbf{L}_{c}$ are first obtained by blending $\textbf{L}_{t,i}$ and $\textbf{L}_{t}$ using a reference image guidance scale $\lambda_{i}$:
+$\textbf{L}_{c}=\frac{\textbf{L}_{t,i}+\lambda_{i}\cdot\textbf{L}_{t}}{1+\lambda_{i}}$ (8)
+Then, the final output logits $\textbf{L}_{\text{final}}$ are computed by combining $\textbf{L}_{c}$ with the fully unconditional logits $\textbf{L}_{\emptyset}$ using the text guidance scale $\lambda_{t}$:
+$\textbf{L}_{\text{final}}=\textbf{L}_{\emptyset}+\lambda_{t}\cdot(\textbf{L}_{c}-\textbf{L}_{\emptyset})$ (9)
+This approach allows flexible control over the influence of both textual and visual conditions during the image editing process.
+In our experiments, for text-to-image generation, we set $\lambda_{t}=20$. For image editing, we set $\lambda_{i}=1$ and $\lambda_{t}=3$.
+Appendix C Additional ablation study
+C.1 Effect on Early and Late Fusion in MLLM.
+Figure 9: Comparison of late fusion and early fusion for multimodal understanding in our setting. We omit the Gen. FFN in this figure for clarity.
+Table 14: Performance comparison of different fusion strategies and distillation methods for multimodal understanding across varying training scales.
+Methods #Trained Tokens of Stage-1 MMB MME-S MMVet SEED AI2D ChartQA TextVQA Avg.
+8B 43.0 1222 13.3 46.8 52.1 10.1 10.6 31.4
+Encoder-based Late Fusion 20B 49.0 1437 16.7 50.4 54.4 11.6 11.3 35.0
+70B 51.7 1426 19.2 55.4 55.9 12.0 19.7 37.8
+8B 42.0 1209 13.2 47.0 53.8 10.5 10.0 31.4
+Decoder-only Early Fusion 20B 45.7 1312 16.7 51.8 56.4 11.7 11.9 34.4
+70B 50.9 1423 16.9 57.4 57.2 14.2 21.0 38.3
+8B 49.4 1327 15.5 56.3 54.4 11.9 11.3 35.3
+Decoder-only Early Fusion 20B 54.3 1410 17.7 61.0 55.4 13.6 15.8 38.3
++ Proposed Distillation 70B 57.6 1476 23.4 63.0 57.2 15.0 25.0 42.0
+300B 60.7 1526 27.1 63.4 60.0 19.2 35.7 45.8
+Figure 10: Performance comparison of different methods for multimodal understanding across varying training scales.
+We conduct an ablation study to evaluate the impact of encoder-based early fusion versus encoder-free late fusion strategies in multimodal understanding, aiming to validate the effectiveness of the decoder-only architecture employed in our model.
+As illustrated in Fig.
+9, late fusion corresponds to the conventional encoder-based MLLM approach, where images are first processed by a vision encoder before being fed into the LLM.
+In contrast, early fusion represents the decoder-only MLLM paradigm of our proposed OneCAT.
+To ensure a fair comparison, the vision encoder (InternViT) in the late fusion model was randomly initialized without pretrained weights, while the LLM in both models was initialized from the pretrained Qwen2.5-1.5B-instruct and kept frozen, aligning with the setup of our stage-1 training.
+All variants are further empoloyed a simplified SFT using LLaVA-665k dataset.
+Fig.
+10 compare the scaling properties of both models with varying trained token budgets for stage-1 and the detailed values are shown in Tab.
+C.1.
+The experimental results demonstrate that the early and late fusion perform on par, while the early fusion model offers a distinct advantage in computational efficiency, aligning with the findings of [82].
+C.2 Effect on Distilling Only Visual Tokens
+We use the same setting of Sec.
+7.5.2 to conduct an ablation study to evaluate the impact of distilling different types of tokens.
+Tab.
+15 shows that distilling only the continuous visual tokens results in a slight overall performance drop, suggesting that it is crucial to distill both visual and text tokens.
+Table 15: Effect of distilling only visual tokens
+Methods MMB MME-S MMVet SEED AI2D ChartQA TextVQA Avg.
+w/o distillation 42.0 1209 13.2 47.0 53.8 10.5 10.0 31.4
+distill only visual tokens 48.1 1299 16.7 55.7 55.2 11.4 10.5 34.8
+[…]
+C.3 Effect on The Increase of Training Tokens of Unified Mid-Training (Stage-2)
+Fig.
+11 provides the performance of our OneCAT-1.5B model on multimodal understanding and generation benchmarks at various checkpoints throughout the unified mid-training stage, corresponding to different amounts of training tokens.
+Before downstream evaluation, the model of each checkpoint undergoes a simplified unified SFT with a combined instruction dataset (LLaVA-665K for understanding and BLIP3o-60K for generation), and we oversample the BLIP3o-60K dataset to achieve a 1:1 training token ratio for understanding and generation tasks.
+Figure 11: Performance of OneCAT-1.5B on different multimodal understanding and generation benchmarks with the increase of training tokens of unified mid-training (Stage-2).
+Appendix D Visualization of discrete visual tokens of different scales
+We generate two $\times 1024$ images and present the visualization of discrete visual tokens across different scales and LLM layers.
+We also visualize the intensity of frequency component by applying Fast Fourier Transform (FFT) to the corresponding tokens’ feature maps.
+As shown in Fig 12, the results show that tokens at lower scales primarily encode low-frequency global information, while higher-scale tokens capture high-frequency details, validating the design rationale of our scale-aware adapter.
+Figure 12: Visualization of discrete visual tokens across scales and LLM layers.
+Left: Each row shows the reshaped feature maps of token hidden states at a specific scale throughout LLM layers.
+The final column displays the feature map of final output codes fed to the image detokenizer for image reconstruction.
+All features are resized to $\times64$ for display.
+Right: Frequency intensity map of the corresponding feature maps.
+Lighter colors indicate larger magnitudes, while pixels closer to the center represent lower frequencies.
+Zoom in better.
+Appendix E More Qualitative results
+In Fig.
+6 and 7, we present qualitative comparisons for text-to-image generation and image editing against several open-source models—including Janus-Pro [15], BAGEL [21], and UniWorld-V1 [57]—as well as the proprietary model GPT-4o-image [72].
+We further present additional qualitative results to comprehensively demonstrate the capabilities of our model.
+Fig.
+13 presents text-to-image generation results from OneCAT under various aspect ratios and resolutions.
+Fig.
+14 showcases OneCAT’s performance on a range of image editing tasks, such as style transfer, object adjustment, attribute modification, object removal, and background editing.
+Additionally, Fig.
+15 provides examples of OneCAT’s multimodal understanding abilities across mathematical reasoning, optical character recognition (OCR), and detailed image captioning.
+Figure 13: Showcase of the text-to-image abilities of the OneCAT model.
+Figure 14: Showcase of the image editing abilities of the OneCAT model.
+Figure 15: Showcase of the visual understanding abilities of the OneCAT model.